### v1

# Autotune: fast, accurate, and automatic tuning parameter selection for LASSO

Abstract
Least absolute shrinkage and selection operator (Lasso), a popular method for high-dimensional regression, is now used widely for estimating high-dimensional time series models such as the vector autoregression (VAR). Selecting its tuning parameter efficiently and accurately remains a challenge, despite the abundance of available methods for doing so. We propose autotune, a strategy for Lasso to automatically tune itself by optimizing a penalized Gaussian log-likelihood alternately over regression coefficients and noise standard deviation. Using extensive simulation experiments on regression and VAR models, we show that $\mathsf{autotune}$ is faster, and provides better generalization and model selection than established alternatives in low signal-to-noise regimes. In the process, $\mathsf{autotune}$ provides a new estimator of noise standard deviation that can be used for high-dimensional inference, and a new visual diagnostic procedure for checking the sparsity assumption on regression coefficients. Finally, we demonstrate the utility of $\mathsf{autotune}$ on a real-world financial data set. An R package based on C++ is also made publicly available on Github.
1 Introduction
In high-dimensional regression with the Lasso, model selection is performed by optimally choosing a single tuning parameter $\lambda$, the penalty level. Recent literature has underscored a need for fast and accurate methods to select $\lambda$ [wu2020survey]. This practice of tuning is also in sharp contrast with the classical practice of regression in low-dimensional settings with a handful of predictors, where one first estimates both the regression coefficients ($\beta$) and the noise standard deviation ($\sigma$). Selection of a predictor in the final model is then governed by a tuning parameter $\alpha$, the significance level, which the practitioner can choose based on their perceived cost of false positive in their context. In this work, we adopt an algorithmic view and develop $\mathsf{autotune}$, a strategy for the Lasso to tune itself by relying only on a pre-specified parameter $\alpha\in(0,1)$. The strategy is faster than existing alternatives, and shows better generalization and model selection in our simulation experiments where signal-to-noise ratio is low.
Formally, for a response vector $\text{Y}=(\text{y}_{i})_{i=1}^{n}\in\real^{n}$ and $p$ predictors in the design matrix $\text{X}=\mathopen{}\mathclose{{\left((\text{x}_{ij})}}\right)_{1\leq i\leq n,1\leq j\leq p}$, Lasso solves the following convex problem
$\min_{\beta\in\real^{p}}\,\mathcal{L}_{\lambda}(\beta):=\,\frac{1}{2n}\|\text{Y}-\text{X}\beta\|^{2}+\lambda\|\beta\|_{1}.$ (2)
The tuning parameter $\lambda$ controls the level of regularization, and should be chosen to strike a balance between false positives and false negatives in model selection. A faster tuning strategy can help build algorithms that repeatedly uses Lasso, e.g. in the analysis of streaming or sequential data in finance [hallac2017network]. A more efficient tuning strategy that performs better model selection with fewer samples can aid new discoveries in large-scale omics problems in biology [ghaemi2019multiomics].
Among the existing tuning strategies, information criteria such as AIC or BIC are fast, but known to perform poorly in high-dimension [wu2020survey, chen2008extended, zhao2015mixture]. This makes cross-validation (CV) the most popular tuning strategy. But even ignoring its ad hoc features (e.g. randomness in data splitting, arbitrary choice of the number of folds), CV is computationally intensive and not preferable in settings that require multiple Lasso fits, e.g. Gaussian graphical models and Vector Autoregressive (VAR) models. CV is also known to overfit [wang2007tuning, lei2020cross] and experience accuracy loss [bates2024cross] in high-dimensional settings. These issues are more prominent for time series data, where a time series cross-validation (TSCV) is used to select folds while respecting the temporal ordering of samples.
To avoid the challenges of tuning, a number of prior works have proposed changing the loss function of Lasso to make it tuning-free, e.g. Scaled Lasso [sun2012scaled], Square Root Lasso [belloni2011square], Rank Lasso [wang2020tuning]. To date, in high-dimensional problems, none of these methods are as fast as the Lasso implemented in glmnet, and their asymptotic theory for time series data is less developed compared to that of Lasso. For VAR models, [kock2025data] has recently proposed an automatic tuning strategy with theoretical guarantees. In this work, instead of changing the loss function, we ask the following question:
Is there a $\lambda$ on the original penalty grid of Lasso which provides a more accurate solution than CV, and if yes, is there a way to find that $\lambda$ quickly?
Figure 1: (1(a)) Mean runtimes (with 1se error bars) of $\mathsf{autotune}$, CV and Scaled Lasso as a function of $p$ increasing from 30 to 1500 with $n=300$ and $\mathopen{}\mathclose{{\left\|{\beta}}}\right\|_{0}=5$. Fig.˜1(b) Mean runtimes of $\mathsf{autotune}$ VAR, TSCV VAR Lasso, and KPS Lasso ˜23 as a function of number of time series (NN) for VAR(1) estimation with $n=150$. Dimension of regression coefficients $p=N^{2}$, so $p$ here varies from 9 to 900. 1(c) Relative test error (metric on y-axis, for details see ˜34) of solution paths of $\mathsf{autotune}$ (blue, whose endpoint is indicated by the blue star symbol) and Lasso(red) against $\log(\lambda)$ for $n=80,p=750$ and $\mathopen{}\mathclose{{\left\|{\beta_{\mathrm{true}}}}}\right\|_{0}=5$. For Lasso, the best $\lambda$ in it’s path is chosen via CV/AIC/BIC, indicated by the labeled vertical lines. For all experiments, we set $\alpha=0.01$ for $\mathsf{autotune}$.
In Fig.˜1, we highlight the performance of $\mathsf{autotune}$ in light of the above question. Fig.˜1(a) shows that autotune is 15-125 times and 52-190 times faster than CV Lasso and Scaled Lasso respectively over a number of simulation scenarios. The speed gains of $\mathsf{autotune}$ become more pronounced in VAR Lasso, where multiple Lasso’s need to be run and tuning is done via TSCV. Fig.˜1(b) shows that $\mathsf{autotune}$ VAR is 73-186 times faster than VAR Lasso tuned via TSCV and 43-97 times faster than the VAR Lasso of [kock2025data] with automatic tuning. In Fig.˜1(c), we illustrate the reason of its speed on a simulated data with $n=80,p=750$, and $\mathopen{}\mathclose{{\left\|{\beta_{\mathrm{true}}}}}\right\|_{0}=5$. Instead of evaluating Lasso solutions on a dense grid of $\lambda$ (red dots), $\mathsf{autotune}$ (blue dots/line) searches over only 3 different data-adaptively chosen $\lambda$ values before stopping (blue star). Interestingly, $\mathsf{autotune}$ picks a $\lambda$ that is nearly optimal for relative test error (RTE, see 34), while AIC, BIC overfit and CV min ˜($\lambda$1) and 1se ˜($\lambda$2) rule both underfit.
[…]
Our $\mathsf{autotune}$ alternately estimates $\beta$ and $\sigma$ in a full negative Gaussian log-likelihood with an $\ell_{1}$ penalty:
$\min_{\beta,\sigma^{2}}\quad\frac{1}{2}\log\sigma^{2}+\frac{1}{2n\sigma^{2}}\|\text{Y}-\text{X}\beta\|^{2}+\lambda_{0}\|\beta\|_{1}.$ (3)
Since for any given $\sigma^{2}$, solving for $\beta$ in (3) is essentially Lasso with tuning parameter $\lambda:=\lambda_{0}\sigma^{2}$, sequentially guessing different values of $\sigma^{2}$ results in fitting Lasso on an automatic, data-driven sequence of $\lambda$ values.
Such an alternating estimation is also at the heart of Scaled Lasso, where $\sigma^{2}$ is estimated using the current $\widehat{\beta}$ as $\widehat{\sigma}^{2}=\|\text{Y}-\text{X}\widehat{\beta}\|^{2}/n$. The automatic tuning of VAR Lasso in [kock2025data] also uses similar updates.
These estimators of $\sigma$ are expected to suffer from the shrinkage bias of $\widehat{\beta}$ throughout the iteration. Our main technical contribution is to use the partial residuals (PR) to correct such shrinkage bias of $\widehat{\sigma}^{2}$ efficiently throughout the iterations.
For a given guess of $\sigma$, and hence a given $\lambda$, the coordinate descent (CD) algorithm for Lasso updates the regression coefficients iteratively using the formulae $\widehat{\beta}_{k}=S_{\lambda}(\sum_{i=1}^{n}\text{x}_{ij}r^{(k)}_{i}/n)$, for $k=1,2,\ldots,p$, where $S_{\lambda}(.)$ is the soft thresholding operator and the coordinates of the partial residuals
$r^{(k)}$ of the $k^{th}$ predictor are defined as
[…]
We observe that for sparse $\beta$, the vector of $p$ standard deviations of PR, $SD(r^{(k)})$, $k=1,2,\ldots,p$, exhibit an interesting behavior throughout the CD iterations. Specifically, if $\beta$ has $s$ non-zero entries, about $s$ many coordinates in the above vector are outliers, i.e. unusually larger than the rest. So, we build a linear model of Y by sequentially introducing these outlier predictors one at a time, using a classical F-test at a significance level $\alpha$ to determine when to stop. The mean squared error (MSE) of this final linear model is used as the new estimator of $\sigma^{2}$. Since we don’t need the multiple linear regression coefficients $\beta$ themselves and only the $\sigma^{2}$, the computation reduces to one simple Gram-Schmidt orthogonalization, which is much faster than fitting $s$ many linear models. Also, in most cases, this estimator $\sigma^{2}$ converges way before the full CD for $\beta$ converges. So, with this new guess of $\sigma^{2}$, autotune moves on to try a new $\lambda$ before waiting for CD to converge in the current $\lambda$. Only when the sequence of $\widehat{\sigma}^{2}$ converges, leading to a final $\lambda$, we compute our final Lasso solution using CD. This offers substantial gain in speed over methods that require first fitting Lasso over a grid of $\lambda$ values.
The $\widehat{\sigma}^{2}$ from $\mathsf{autotune}$ shows promising results in high-dimensional inference compared to alternatives such as Scaled Lasso. By tracking the gain in R2 in the nested linear models above, we can also design a visual diagnostic analogous to the scree plot in PCA. To the best of our knowledge, this is the first diagnostic tool for checking the sparsity assumption in Lasso.
In a wide range of simulation experiments using data generating process from [hastie2020best], we benchmark $\mathsf{autotune}$ against CV-tuned and Scaled Lasso. In Sec.˜4.1.2, $\mathsf{autotune}$ shows higher accuracy than both CV-tuned Lasso and Scaled Lasso in terms of estimation (Relative MSE), prediction (“Relative Test Error” [Sec. 3.1, hastie2020best]), and variable selection (Area under ROC curve). In Sec.˜4.1.3, $\mathsf{autotune}$ demonstrates 15-200 times speedup over CV-tuned Lasso and Scaled Lasso. In our simulations for VAR models in Sec.˜4.2, we find that accuracy gains of $\mathsf{autotune}$ over benchmarks on the same metrics are even larger, and $\mathsf{autotune}$ is even faster over TSCV-tuned VAR Lasso and [kock2025data]. Lastly, we show that $\mathsf{autotune}$ improves $\sigma^{2}$ estimation in high-dimensional data as compared to the state-of-the-art methods, leading to better high dimensional inference [dezeure2015high] in Sec.˜4.3.1 and provides promising diagnostics for checking the sparsity assumption in Sec.˜4.3.2.
We also perform experiments on a real-life dataset sp500 (given in Scaled Lasso’s R package, scalreg [scalreg]) in Sec.˜5, and demonstrate the gains of $\mathsf{autotune}$ Lasso over the benchmarks (including Scaled Lasso) in terms of relative test error, Out-of-sample R2 and variable selection. The other sections of the paper are organized as follows. Sec.˜2 explains the prior state-of-the-art algorithms for computing Lasso and tuning $\lambda$(most of which are used as benchmarks for measuring $\mathsf{autotune}$’s performance). Sec.˜3 elucidates the algorithm of $\mathsf{autotune}$ Lasso, $\mathsf{active}$ $\mathsf{autotune}$ Lasso, and $\mathsf{autotune}$ VAR.
We also provide the R package (https://github.com/Tathagata-S/Autotune) for a fast implementation of $\mathsf{autotune}$ Lasso, whose inner layers are coded in C++. All our simulations in this paper are run in R [Rcoreteam].
2 Background
In this section, we describe existing tuning strategies for Lasso in regression and VAR modeling, and discuss their limitations. Then we briefly describe partial residuals (PR) in CD algorithms for Lasso, and sequential F-tests for model selection in classical statistics - two key technical ingredients used in $\mathsf{autotune}$.
Tuning with cross-validation (CV):
Tuning $\lambda$ via K-fold cross-validation [hastie2015statistical] remains the most common tuning strategy among practitioners. Denote $\widehat{\beta}\mathopen{}\mathclose{{\left(\lambda}}\right)=\mathop{\mathrm{argmin}}_{\beta\in\real^{p}}\mathcal{L}_{\lambda}(\beta)$, (see ˜2). Once we have $\widehat{\beta}(\lambda)$ on a grid of $\lambda$ values, there are two choices
  1. ($\lambda$1)
[…]
  2. ($\lambda$2)
    CV(1se): largest $\lambda$ such that its K - fold CV error was within 1 SD of the minimum CV error.
Despite its popularity, the number of folds is set to 5 or 10 in an ad hoc manner. The final chosen $\lambda$ ˜($\lambda$1) is also known to be susceptible to the fold assignment used in CV [bovelstad2007predicting, roberts2014stabilizing]. Recent works have also shown that CV overfits the data [wang2007tuning, lei2020cross] and loses accuracy [bates2024cross]. To the best of our knowledge, there has been no systematic study to assess the quality of CV(1se) ˜($\lambda$2) solutions.
Information Criterion (IC) based tuning:
AIC [akaike2003new] and BIC [schwarz1978estimating] introduced the idea of penalizing complexity of the model fitted for inducing model selection. The general formula is $n\log\mathopen{}\mathclose{{\left\|{\text{Y}-\text{X}\widehat{\beta}}}}\right\|_{2}^{2}+kp$ where $k=2$ for AIC and $k=\log(n)$ for BIC. For tuning $\lambda$ in Lasso, AIC and BIC’s simplicity, ease of computation, and avoiding repeated data-splitting calculations made them a formidable and popular alternative to CV. But these information criteria were formulated under the assumption that $p$ is fixed and sample size $n\to\infty$, making them unsuitable for high dimensional regime $n<p$ [chen2008extended, zhao2015mixture]. Other information criterion like eBIC [chen2008extended], HBIC [zhao2015mixture] are also proposed to combat scenarios with $p$ growing with $n$ but are not used widely by practitioners and recent research [10.1093/jjfinec/nbab023] suggests they don’t improve upon the performance of BIC in time series analysis.
Linking $\lambda$ with error variance
High-dimensional asymptotics of Lasso for linear model with Gaussian errors suggests choosing $\lambda$ in the order of $\sigma\sqrt{\log(p)/n}$ [bickel2009simultaneous]. Since $\sigma$ is unknown in practice, a line of work reformulate the loss function ˜2 or ˜3 to make the tuning of $\lambda$ independent of $\sigma$ [belloni2011square, lederer2015don, wang2020tuning].
Another line of work, related more to our method, focuses on estimating $\sigma$ from data and using it for tuning $\lambda$. For example, Scaled Lasso [sun2012scaled] introduced a $\sigma$-based tuning procedure for $\lambda$ by modifying the penalized Gaussian log-likelihood objective ˜3.
$\min_{\beta,\sigma^{2}}\quad\sigma/2+\frac{1}{2n\sigma}\|\text{Y}-\text{X}\beta\|^{2}+\lambda_{0}\|\beta\|_{1}.$ (5)
˜5 was motivated by the discussions [sun2010comments, antoniadis2010comments] on the paper [stadler2010]. Scaled Lasso [sun2012scaled] proposes an iterative minimization algorithm alternating over $\beta$ and $\sigma^{2}$ for tuning $\lambda$ in Lasso: use the $\widehat{\beta}$’s from the Lasso solution path to calculate the RSS for estimating the $\sigma$, and then the $\widehat{\sigma}$ tunes the $\lambda$ for the next iteration of coordinate descent. Moreover, [sun2012scaled] derived asymptotic guarantees of Scaled Lasso for estimating $\sigma$.
Additionally, [yu2019estimating] proposed Natural and Organic Lasso which optimized the following respective modified loss functions
$\min_{\frac{1}{\sigma^{2}},\frac{\beta}{\sigma^{2}}\in\real_{+}\times\real^{p}}$ $\,\,-\frac{1}{2}\log\mathopen{}\mathclose{{\left(\frac{1}{\sigma^{2}}}}\right)+\frac{1}{2n\sigma^{2}}\mathopen{}\mathclose{{\left\|{\text{Y}-\text{X}\frac{\beta/\sigma^{2}}{1/\sigma^{2}}}}}\right\|_{2}^{2}+\lambda\|\beta/\sigma^{2}\|_{1},\quad\text{and}\quad\min_{\beta\in\real^{p}}\,\,\frac{1}{2n}\|\text{Y}-\text{X}\beta\|^{2}+\lambda\|\beta\|_{1}^{2}$ (6)
[yu2019estimating] arrived at Natural Lasso by making the original Lasso objective ˜2 scale invariant. Natural Lasso still needs CV to tune its $\lambda$.
But it’s known that $\widehat{\beta}$ from Lasso’s coordinate descent has shrinkage bias [hastie2015statistical] and using its corresponding RSS will lead to $\widehat{\sigma}$ with large finite-sample errors. Our simulations(Sec.˜4.3) report the same finite sample difficulties after running an extensive simulation study on $\sigma$ estimation in Lasso regression. Finite sample limitations of Scaled Lasso was also reported in [yu2019estimating]’s simulations.
Tuning $\lambda$ in VAR models:
Vector Autoregression (VAR) is a workhorse model in multivariate time series data, widely used in macroeconomics [banbura2010large, bernanke2005measuring], portfolio selection in finance [fan2011sparse], identifying gene regulatory networks in genomics [michailidis2013autoregressive, broll2025prolong] and for understanding functional connectivity between different brain regions in neuroscience [smith2012future, seth2013granger]. Formally, a VAR($d$) model on a $p$-dimensional stationary time series $\mathopen{}\mathclose{{\left\{Z^{t}}}\right\}_{t\geq 1}=\mathopen{}\mathclose{{\left\{\mathopen{}\mathclose{{\left(Z^{t}_{1},\dots,Z^{t}_{p}}}\right)}}\right\}_{t\geq 1}$ with no contemporaneous association takes the form
$\mathrm{VAR}\mathopen{}\mathclose{{\left(d}}\right):Z^{t}=\sum_{i=1}^{d}A_{i}Z^{t-i}+\epsilon^{t},\quad\epsilon^{t}\stackrel{{\scriptstyle\textrm{i.i.d.}}}{{\sim}}\mathopen{}\mathclose{{\left(0,\Sigma_{\epsilon}}}\right),\quad\Sigma_{\epsilon}=\mathop{\mathrm{diag}}\mathopen{}\mathclose{{\left(\sigma^{2}_{1},\dots,\sigma^{2}_{p}}}\right),$ (7)
where $A_{1},\dots,A_{d}$ are $p\times p$ transition matrices which need to be estimated. Given $n+d$ stationary observations of $\mathopen{}\mathclose{{\left\{Z^{t}}}\right\}_{t\geq 1}$, all the $d$ VAR transition matrices can be estimated in the autoregressive design [hyndman2018forecasting]
$\underbrace{\mathopen{}\mathclose{{\left[\begin{array}[]{c}\mathopen{}\mathclose{{\left(Z^{n+d}}}\right)^{\top}\\ \vdots\\ \mathopen{}\mathclose{{\left(Z^{d+1}}}\right)^{\top}\end{array}}}\right]}_{\text{Y}_{n\times p}}=\underbrace{\mathopen{}\mathclose{{\left[\begin{array}[]{cccc}\mathopen{}\mathclose{{\left(Z^{n+d-1}}}\right)^{\top}&\mathopen{}\mathclose{{\left(Z^{n+d-2}}}\right)^{\top}&\dots&\mathopen{}\mathclose{{\left(Z^{n}}}\right)^{\top}\\ \vdots&\vdots&\vdots&\vdots\\ \mathopen{}\mathclose{{\left(Z^{d}}}\right)^{\top}&\mathopen{}\mathclose{{\left(Z^{d-1}}}\right)^{\top}&\dots&\mathopen{}\mathclose{{\left(Z^{1}}}\right)^{\top}\end{array}}}\right]}_{\text{X}_{n\times dp}}\underbrace{\mathopen{}\mathclose{{\left[\begin{array}[]{c}A_{1}^{\top}\\ A_{2}^{\top}\\ \vdots\\ A_{d}^{\top}\end{array}}}\right]}_{\Phi_{dp\times p}}+\underbrace{\mathopen{}\mathclose{{\left[\begin{array}[]{c}\mathopen{}\mathclose{{\left(\varepsilon_{n+d}}}\right)^{\top}\\ \vdots\\ \mathopen{}\mathclose{{\left(\varepsilon^{d+1}}}\right)^{\top}\end{array}}}\right]}_{E_{n\times p}},$ (21)
where $\Phi$ is the $dp\times p$ matrix constructed by stacking $A_{1}^{\top},A_{2}^{\top},\dots,A_{d}^{\top}$. The parameter space dimension ($dp^{2}$) grows quadratically with the number of time series $p$, and the OLS estimate of $\Phi$ will be consistent only if $n>dp^{2}$, which is often a big ask for sample size in modern VAR applications. So, it is common to assume $\Phi$ is sparse and use $\ell_{1}$-penalized least squares regressions [basu2015regularized]. We obtain the VAR Lasso estimates of $\Phi$ by conducting $p$ separate, columnwise Lasso regressions. Mathematically,
$\widehat{\Phi}_{\cdot,i}=\underset{\beta\in\mathbb{R}^{dp}}{\operatorname{argmin}}\frac{1}{2n}\mathopen{}\mathclose{{\left\|\text{Y}_{\cdot,i}-\text{X}\beta}}\right\|^{2}+\lambda_{i}\mathopen{}\mathclose{{\left\|{\beta}}}\right\|_{1},\quad\text{ for }i=1,\ldots,p.$ (22)
Time Series Cross-validation (TSCV):
A common approach to tune these $p$ different $\lambda_{i}$’s is the time series CV (TSCV) [Sec. 5.10, hyndman2018forecasting]. It is more constrained because the temporal order of the data must be preserved when choosing the different folds. A “rolling window” is adopted, where the dataset is divided into $2K$ folds along the temporal dimension. Then the $(K+1)^{th}$ fold is assigned the role of a test set, and all the folds corresponding to previous time points are used as the training data to fit the VAR Lasso on the original grid of $\lambda_{i}$s. We repeat this process $K$ times, where every fold in the latter half plays the role of test fold once, and in each repetition, all the previous folds are used as training data. Lastly, TSCV averages the $K$ mean squared errors (MSE) for each $\lambda_{i}$ to get the mean TSCV error, and chooses a $\lambda_{i}$ which minimizes this. As a result, TSCV has some statistical limitations over the usual CV [bergmeir2012use, bergmeir2018note]. For example, constraints in data-splitting for TSCV reduces the effective sample size available for tuning of $\lambda$.
Moreover, running separate TSCV on each column can be computationally prohibitive for analysis of modern day large time series datasets. Hence, prior software [bigtimeRpackage] have often focused on uniform penalization across all the separate column-wise lasso regressions, which is inadequate for handling heterogeneity across multiple time series.
Automatic Tuning for VAR Lasso:
Recently [kock2025data] proposed an automated tuning mechanism for VAR Lasso, building upon the iterative algorithm proposed in [belloni2012sparse]. Denote $\widehat{\Phi}^{(t)}$ as the estimate of $\Phi$ at the $t^{th}$ iteration. Then, [kock2025data] iteratively computes the following columnwise weighted lasso estimator for $t=1,2,3,\ldots$
$\widehat{\Phi}^{(t)}_{\cdot,i}=\underset{\beta\in\mathbb{R}^{dp}}{\operatorname{argmin}}\frac{1}{n}\mathopen{}\mathclose{{\left\|\text{Y}_{\cdot,i}-\text{X}\beta}}\right\|^{2}+\frac{\lambda}{n}\mathopen{}\mathclose{{\left\|{\Gamma^{(t)}_{i}\beta}}}\right\|_{1},\quad\text{ for }i=1,\ldots,p$ (23)
$\text{where }\Gamma^{(t)}_{i}=\mathop{\mathrm{diag}}\mathopen{}\mathclose{{\left(\widehat{v}^{(t)}_{i,1},\dots,\widehat{v}^{(t)}_{i,pq}}}\right),\quad\text{ with }\widehat{v}^{(t)}_{i,j}=\frac{1}{\sqrt{n}}\sqrt{\sum_{r=1}^{n}\text{X}_{r,j}^{2}\mathopen{}\mathclose{{\left(\text{Y}_{r,i}-\text{X}_{r:}\widehat{\Phi}^{(t-1)}_{\cdot,i}}}\right)^{2}}.$ (24)
Here the weights $\Gamma_{i}$ are adjusted according to the residuals in the previous iteration. This strategy is similar to the iterative algorithm precursor to scaled lasso [wu2020survey], and shares its limitation of shrinkage bias in $\widehat{\Phi}$ that are used in noise estimator $\widehat{v}$. In Sec.˜4.2 we compare different tuning strategies in VAR Lasso including $\mathsf{autotune}$, TSCV, AIC, BIC, and [kock2025data] on simulated time series data.
Coordinate descent(CD) and partial residuals:
Coordinate descent’s ease of implementation, availability of coordinate-wise closed-form updates, augmented with all the computational tricks of warm starts, and active set selection makes it the default and the fastest algorithm [10.1214/07-AOAS131] for researchers using Lasso over other algorithms like proximal gradient descent [nesterov2013introductory], and LARS [efron2004least]. Additionally, it enjoys convergence guarantees [tseng2001convergence]. CD’s idea is first to fix the $\lambda$ in ˜2 and then iteratively optimize over each parameter holding other parameters constant at their current values.
Input: Y, Normalized predictors X (i.e., $\mathopen{}\mathclose{{\left\|{\text{X}_{\cdot,j}}}}\right\|_{2}^{2}=n$), tuning parameter $\lambda$
1 Initialize $\widehat{\beta}^{(0)}=0_{p}$, $r=$ $\text{X}\widehat{\beta}^{(0)}=Y$
2 for $t=1,\ldots,\mathsf{max.iterations}$ do
3  for $j=1,\ldots,p$ do
4     $\widehat{\beta}^{(t)}_{j}=\mathcal{S}_{\lambda}\mathopen{}\mathclose{{\left(\frac{1}{n}\text{X}_{\cdot,j}^{\top}r^{(j)}}}\right)=\mathcal{S}_{\lambda}\mathopen{}\mathclose{{\left(\frac{1}{n}\text{X}_{\cdot,j}^{\top}\mathopen{}\mathclose{{\left(r+\text{X}_{\cdot,j}\widehat{\beta}^{(t-1)}_{j}}}\right)}}\right)$
5     $r\leftarrow r-\text{X}_{\cdot,j}\mathopen{}\mathclose{{\left(\widehat{\beta}^{(t)}_{j}-\widehat{\beta}^{(t-1)}_{j}}}\right)$
6  end for
7 If $\mathopen{}\mathclose{{\left\|{\widehat{\beta}^{(t)}-\widehat{\beta}^{(t-1)}}}}\right\|$ is small: Break
8 end for
Output: $\widehat{\beta}^{(t)}$
Algorithm 1 Pseudocode of Coordinate Descent for Lasso
We provide the pseudocode of Lasso at a fixed $\lambda$ in Algorithm 1. A core component of CD updates in Lasso is the partial residual $r^{(j)}$, $\forall j\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$, which comes out to be $r^{(j)}=r+\text{X}_{.,j}^{\top}\widehat{\beta}_{j}$ (from ˜4), where $r=\text{Y}-\sum_{k\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}}\text{X}_{.,k}^{\top}\widehat{\beta}_{k}$ is the full residual. Partial residuals are continuously computed and updated inside every CD step, but there has been only a limited discussion about it in the prior literature. We take a deeper look into the partial residuals in the next section.
[…]
Sometimes, to further reduce computational costs, a subset of coefficients called “active set” is selected via some tests [10.1111/j.1467-9868.2011.01004.x, hastie2015statistical] at each $\lambda$ in the grid to gather all the predictors whose absolute correlation with the YY/residuals is not below a certain threshold depending on $\lambda$. This ensures that the predictors will contribute to the prediction of $Y$ at that $\lambda$. Coefficients outside the active set have the value 0, and the CD algorithm is iterated only over the active set.
In the state-of-the-art algorithms, scaled Lasso internally uses LARS, which is computationally slower than CD [10.1214/07-AOAS131], and K-fold CV Lasso computes the entire Lasso solution path $K+1$ times. autotune Lasso runs the coordinate descent(CD) once at each $\lambda$ in its smaller regularization path(than CV lasso, see Fig.˜4). Therefore, we may therefore expect $\mathsf{autotune}$ to be faster (Sec.˜4.1.3, 4.2).
Sequential F-tests.
Sequential F-tests/sum of squares [Chap 4.6, montgomery2020introduction] is a classic model selection tool which is often employed to calculate ANOVA tables [carpackage]. Given the responses Y and $p$ predictors Xj, $j\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$, it looks into the following sequence of nested linear models and their corresponding residual sum of squares (RSS):
$\mathcal{M}_{k}:$ $\text{ Y}=\text{ X}_{1}\beta_{1}+\text{ X}_{2}\beta_{2}+\dots+\text{X}_{k}\beta_{k}+\varepsilon,\qquad k\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$ (25)
$\text{RSS}(\mathcal{M}_{k})$ $=\sum_{i=1}^{n}\mathopen{}\mathclose{{\left(\text{y}_{i}-\widehat{\text{y}}_{i}}}\right)^{2}\quad\text{where $\widehat{\text{y}}_{i}$'s comes from fitting $\mathcal{M}_{k}$ to data.}\quad$ (26)
Then, a sequential F-test iteratively performs the hypothesis test $\mathcal{H}_{0,k}:\mathopen{}\mathclose{{\left\{\beta_{k}=0}}\right\}|\mathopen{}\mathclose{{\left(\beta_{1},\dots,\beta_{k-1}}}\right)$ using the statistic
$F=\frac{\text{RSS}(\mathcal{M}_{k-1})-\text{RSS}(\mathcal{M}_{k})}{\text{RSS}(\mathcal{M}_{k})/(n-k)}\underset{\mathcal{H}_{0,k}}{\sim}F_{1,n-k}$ (27)
until either we arrive at a $k_{0}\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$ where we fail to reject the $\mathcal{H}_{0,k_{0}}$ or we reject $\mathcal{H}_{0,j}$ $\forall j\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$. If the procedure stops, we conclude that the significant predictors for the regression of Y on X are $\mathopen{}\mathclose{{\left\{1,\dots,k_{0}-1}}\right\}$, otherwise none of the predictors are spurious. Note that in $n<p$ regime, we will always arrive at a $k_{0}\leq n$ where the sequential F-test stops. The rank order of predictors in which the nested models $\mathcal{M}_{k}$ are constructed is important for selecting the final model.
3 Method: tuning with $\mathsf{autotune}$
Consider the negative penalized Gaussian log likelihood ˜28, a biconvex function in $(\beta,\sigma^{-2})$:
$\min_{\beta,\sigma^{-2}}\quad$ $f(\beta,\sigma^{-2}):=-\frac{1}{2}\log\sigma^{-2}+\frac{\sigma^{-2}}{2n}\|\text{Y}-\text{X}\beta\|^{2}+\lambda_{0}\|\beta\|_{1}.$ (28)
Note that any alternating estimation method of the following form can lead to an automatic tuning procedure for Lasso with a data-driven sequence of tuning parameters $\lambda_{n}=\lambda_{0}\,\widehat{\sigma}^{2}_{n}$, for $n\geq 1$.
  1. (M1)
    Given $\sigma^{-2}=\widehat{\sigma}^{-2}$, update $\widehat{\beta}$ by running Algorithm 1 with $\lambda=\lambda_{0}\widehat{\sigma}^{2}$,
  2. (M2)
    Given $\beta=\widehat{\beta}$, update $\widehat{\sigma}^{-2}$ using the outputs of Algorithm 1.
In fact, the original version of Scaled Lasso algorithm proposed in [sun2010comments] falls in this general framework, with $\widehat{\sigma}^{2}$ in ˜(M2) being estimated by the conditional maximum likelihood estimator $\|\text{Y}-\text{X}\widehat{\beta}\|^{2}/n$. The KPS algorithm for VAR Lasso [kock2025data] can also be viewed as the same strategy employed on $p$ different noise variance parameters. A potential shortcoming of this approach is that it does not regularize $\sigma^{2}$ in any way. Hence in the Scaled Lasso paper [sun2012scaled], the authors proposed a form of degree of freedom correction $\widehat{\sigma}^{2}=\|\text{Y}-\text{X}\widehat{\beta}\|^{2}/(n-a)$, for some $a>0$.
In our empirical analyses, we found that these strategies of iteratively updating $\sigma$ suffer from the shrinkage bias in $\widehat{\beta}$. So we develop a more aggressive bias correction strategy based on the partial residuals (PR) which are already being computed inside Algorithm 1.
A basic autotune strategy.
Assuming the true $\beta$ is $s$-sparse and $s\ll p$, we expect that the standard deviations (SD) of the PR corresponding to some $s$ predictors will be somewhat larger than the standard deviation of the other (ps)(p-s) ones. Betting on this general premise, we rank the SD of the $p$ PRs at the end of Algorithm 1 from largest to smallest, and use this rank list to sequentially add predictors to build a nested sequence of linear models using F-tests at the significance level $\alpha$ (see Background Sec.˜2). The MSE of this final linear model is then taken as the new estimator of $\widehat{\sigma}^{2}$.
Since for conducting these F-tests, we only require the RSS and not all the multiple linear regression coefficients from these linear models, we can reduce the computation substantially by performing a single Gram-Schmidt (GS) orthogonalization of $\mathopen{}\mathclose{{\left\{\text{X}_{\cdot,1},\text{X}_{\cdot,2},\dots}}\right\}$. Without loss of generality, let the rank list of predictors be $\{1,2,\ldots,p\}$. Then we define $\mathopen{}\mathclose{{\left\{u_{\cdot,1},\dots,u_{\cdot,k-1}}}\right\}$ to be the set of first k1k-1 predictors in the rank list orthogonalized via GS, and $\epsilon^{(k-1)}$ to be the vector of residuals in the linear model  $\text{ Y}\sim u_{\cdot,1}+u_{\cdot,2}+\dots+u_{\cdot,k-1}$. Formally,
$u_{\cdot,k}$ $=\text{X}_{\cdot,k}-\sum_{i=1}^{k-1}\frac{\langle\text{X}_{\cdot,k},u_{\cdot,i}\rangle}{\langle u_{\cdot,i},u_{\cdot,i}\rangle}u_{\cdot,i},\quad\epsilon^{(k)}=\epsilon^{(k-1)}-\frac{\langle{\epsilon^{(k-1)}},{u_{\cdot,k}}\rangle}{\langle{u_{\cdot,k}},{u_{\cdot,k}}\rangle}u_{\cdot,k}.$ (29)
Note that $\epsilon^{(k)}$ is essentially the vector of residuals from the linear model $\mathcal{M}_{k-1}$ of $Y$ on $\mathopen{}\mathclose{{\left\{\text{X}_{\cdot,1},\text{X}_{\cdot,2},\dots}}\right\}$. Then, we can perform F-test of $\mathcal{M}_{k-1}$ against $\mathcal{M}_{k}$ (the linear model we get after adding $u_{\cdot,k}$ to $\mathcal{M}_{k-1}$) as follows:
$\mathrm{RSS}(\mathcal{M}_{k})$ $=\mathopen{}\mathclose{{\left\|{\epsilon^{(k)}}}}\right\|_{2}^{2}\quad\text{and}\quad\mathrm{RSS}(\mathcal{M}_{k-1})=\mathopen{}\mathclose{{\left\|{\epsilon^{(k-1)}}}}\right\|_{2}^{2}$ (30)
and we use the formula in ˜27. Let $k_{0}-1$ be the number of predictors in the final linear model. This procedure reduces the computational complexity to $\mathcal{O}(n.k_{0})$ from $\mathcal{O}(n.k_{0}^{2})$ in the basic procedure where each model $\mathcal{M}_{k}$ is fully fitted without using the previous fit $\mathcal{M}_{k-1}$. The full details of this $\sigma$ update is given in Algorithm 2. So, this version of $\mathsf{autotune}$ is effectively using Algorithms 1 and 2 alternately to perform steps ˜(M1) and ˜(M2) above.
Early stopping in Algorithm 1.
We observed that $\widehat{\sigma}^{2}$, which is a uni-dimensional nuisance parameter, converged way faster as compared to the high-dimensional $\widehat{\beta}$ in Algorithm 1 (demonstrated in Fig.˜3). Hence, we set an additional criterion for flagging the quick convergence of $\sigma$ estimates. The $\mathsf{autotune}$ stops updating $\widehat{\sigma}$ at the iteration when the estimated support set is a subset of that in the previous iteration. Then fixing the final $\widehat{\sigma}$ (hence fixing $\lambda=\lambda_{0}\widehat{\sigma}^{2}$), $\mathsf{autotune}$ keeps on repeating ˜(M1) till the $\widehat{\beta}$’s converge. The final version is described in Algorithm 3.
Input: Y, X, $\alpha$, $r$, $\widehat{\beta}$
1
21exCalculate partial residuals $r^{(j)}\leftarrow r+\text{X}_{\cdot,j}\widehat{\beta}_{j}$ $\forall j\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$
3
[…]
5
[…]
7
[…]
9[2pt] for $i=1,\ldots,p$ do
10  $j\leftarrow\mathsf{Predictor.Ranking}[i]$
11 [2pt] if $i>1$ then
12     $u_{\cdot,i}=\text{X}_{\cdot,j}-\sum_{k=1}^{i-1}\frac{\langle\text{X}_{\cdot,j},u_{\cdot,k}\rangle}{\langle u_{\cdot,k},u_{\cdot,k}\rangle}u_{\cdot,k}$   # Gram-Schmidt Process
13  end if
[…]
15 [2pt] Perform Sequential F-test ˜27 using F-test statistic = $\frac{\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\|{\text{Y}^{\mathsf{(temp)}}}}}\right\|_{2}^{2}-\mathrm{RSS}(\mathcal{M}_{i})}}\right)}{\mathrm{RSS}(\mathcal{M}_{i})/(n-i)}$, and cutoff = $F_{\alpha;1,n-i}$
16 [2pt] if $\mathrm{F-test}$ $\mathrm{statistic}\leq\mathrm{cutoff}$ then
17     break and go to line 18.
18  end if
19 Include $j^{th}$ predictor $\text{X}_{\cdot,j}$ in the $\mathsf{Support.Set}$
20 [2pt] $\text{Y}^{\mathsf{(temp)}}\leftarrow\text{Y}^{\mathsf{(temp)}}-\widehat{\text{Y}}$
21 end for
$2\widehat{\sigma}^{2}\leftarrow\frac{1}{n-\mathopen{}\mathclose{{\left|\mathsf{Support.Set}}}\right|}\sum_{i=1}^{n}\mathopen{}\mathclose{{\left(\text{y}_{i}^{\mathsf{(temp)}}}}\right)^{2}$
Output: new noise variance estimate $\widehat{\sigma}^{2}$, updated support set $\mathsf{Support.Set}$, $\mathsf{Predictor.Ranking}$.
Algorithm 2 Sigma Update – Updates the estimate of noise variance $\widehat{\sigma}^{2}$.
Input: Y, Normalized predictors X (i.e., $\mathopen{}\mathclose{{\left\|{\text{X}_{\cdot,j}}}}\right\|_{2}^{2}=n$), Sequential F-test significance level $\alpha$
1
21exInitialize $\widehat{\beta}\leftarrow 0$, $r\leftarrow\text{Y}$, $\widehat{\sigma}^{2}\leftarrow\mathrm{Var}(\text{Y})$, $\lambda_{0}\leftarrow\frac{1}{\mathrm{Var}(\text{Y})}\|\frac{\text{X}^{\top}\text{Y}}{2n}\|_{\infty}$, $\mathsf{Predictor.Ranking}\leftarrow\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$, $\mathsf{Support.Set}\leftarrow\phi$, sigma.update.flag $\leftarrow\mathsf{TRUE}$
3
41exwhile $\mathrm{error}\geq\mathsf{error.tolerance}$ do
5  Set $\lambda=\lambda_{0}\widehat{\sigma}^{2}$
6  Set $\mathsf{Support.Set}^{\mathsf{(old)}}\leftarrow\mathsf{Support.Set}$, $\widehat{\beta}^{\mathsf{(old)}}\leftarrow\widehat{\beta}$
7 [2pt] for $i=1,\ldots,p$ do
8     $j\leftarrow\mathsf{Predictor.Ranking}[i]$
9    [2pt] $\widehat{\beta}_{j}\leftarrow\mathsf{Soft.Threshold}_{\lambda}\mathopen{}\mathclose{{\left(\frac{1}{n}\text{X}_{\cdot,j}^{\top}r+\widehat{\beta}_{j}}}\right)$
10    [2pt] $r\leftarrow r+\text{X}_{\cdot,j}\mathopen{}\mathclose{{\left(\widehat{\beta}^{\mathsf{(old)}}_{j}-\widehat{\beta}_{j}}}\right)$
11  end for
12
13 if $\mathrm{sigma.update.flag}==\mathsf{TRUE}$ then
[…]
15     if $\mathsf{Support.Set}\subseteq\mathsf{Support.Set}^{\mathsf{(old)}}$ then
16        sigma.update.flag $\leftarrow\mathsf{FALSE}$
17
18     end if
19
20  end if
21 $\mathrm{error}\leftarrow\|\widehat{\beta}-\widehat{\beta}^{\mathsf{(old)}}\|_{1}\bigg/\|\widehat{\beta}^{\mathsf{(old)}}\|_{1}$
22 end while
Output: regression coefficients’ estimate $\widehat{\beta}$, noise variance estimate $\widehat{\sigma}^{2}$, regularization parameter $\lambda$
Algorithm 3 Autotune Lasso – Return estimates of regression coefficients $\beta$.
Warm starts and initialization of $\lambda_{0}$.
We use a warm start strategy as is common in glmnet. This means we initialize the algorithm for $\lambda_{k+1}$ at the solution obtained at the previous candidate value of $\lambda_{k}$, for every $k=1,2,\ldots$. Even though our algorithm searches over a sequence of $\lambda$ values which are not necessarily close, we find that this strategy helps with quick convergence of CD as opposed to fixed initializers such as 0.
Asymptotic theory suggests $\lambda\asymp\mathopen{}\mathclose{{\left\|\text{X}^{\top}\varepsilon/n}}\right\|_{\infty}\asymp\sigma\sqrt{\log p/n}$ for Lasso. Adjusting for the loss function in $\mathsf{autotune}$, $\lambda_{0}\asymp\mathopen{}\mathclose{{\left\|\text{X}^{\top}\varepsilon/n}}\right\|_{\infty}/\text{Var}(\varepsilon)\asymp\sigma\sqrt{\log p/n}/\sigma^{2}$ is a natural choice. Since $\varepsilon$ is unknown in practice, a reasonable starting point for our iterative algorithm is to use $Y$ instead of $\varepsilon$ in the above expression.
Prior literature [glmnet, hastie2015statistical] recommends initiating computation of Lasso path at $\lambda=\mathopen{}\mathclose{{\left\|{\text{X}^{\top}\text{Y}/n}}}\right\|_{\infty}$ which is just big enough so that $\widehat{\beta}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\|{\text{X}^{\top}\text{Y}/n}}}\right\|_{\infty}}}\right)=0$ (recall $\widehat{\beta}\mathopen{}\mathclose{{\left(\lambda}}\right)=\mathop{\mathrm{argmin}}_{\beta\in\real^{p}}\mathcal{L}_{\lambda}(\beta)$). But in $\mathsf{autotune}$, we found out this approach makes the regression estimates $\widehat{\beta}$ stuck at 0. We started seeing the empirical success of $\mathsf{autotune}$ after beginning the Lasso path at a $\lambda$ slightly smaller than $\mathopen{}\mathclose{{\left\|{\text{X}^{\top}\text{Y}/n}}}\right\|_{\infty}$. Moreover, as $\mathsf{autotune}$ starts from $\widehat{\beta}=0$, we set $\widehat{\sigma}^{2}_{\mathsf{(initial)}}=$Var(Y). Consequently, our final choice of $\lambda_{0}$ ensures that the initial $\lambda=\lambda_{0}\widehat{\sigma}^{2}_{\mathsf{(initial)}}=\frac{1}{2}\mathopen{}\mathclose{{\left\|{\text{X}^{\top}\text{Y}/n}}}\right\|_{\infty}$.
Speeding up predictor ranking.
Switching from standard deviation, which is essentially the $\ell_{2}$ norm of PR, to an $\ell_{1}$-norm of PR helps improve speed without comprising accuracy. In other words, important predictors are still ranked high in terms of both norms.
Active set selection.
We can also utilize active set selection [10.1111/j.1467-9868.2011.01004.x] to reduce the computational load further. We know from the Lasso’s KKT conditions [hastie2015statistical] that for all predictors $\text{X}_{\cdot,j}$ with nonzero estimates of coefficients, $\mathopen{}\mathclose{{\left|\langle{\text{X}_{\cdot,j}},{\text{Y}-\text{X}\widehat{\beta}\mathopen{}\mathclose{{\left(\lambda}}\right)}\rangle/n}}\right|=\lambda$ and the remaining predictors will have smaller absolute correlation. A reasonable expectation is that at the start of CD at a $\lambda$, predictors with a small inner product with the residuals are more likely to have a zero coefficient as compared to the predictors with bigger inner products. We use this intuition to remove the spurious predictors, which can potentially boost the speed of the computations while still arriving at the optimal minimizer (details in Active-Autotune-Lasso, and its empirical performance is demonstrated in Sec.˜A.2).
3.1 Autotune for VAR Lasso
Estimating a VAR($d$) model on a $p$-dimensional time series under $\ell_{1}$ penalty is equivalent to solving $p$ separate lasso regression problems (Sec.˜2, ˜22). For a $p$-dimensional stationary time series $\mathopen{}\mathclose{{\left\{Z^{t}}}\right\}_{t\geq 1}=\mathopen{}\mathclose{{\left\{\mathopen{}\mathclose{{\left(Z^{t}_{1},\dots,Z^{t}_{p}}}\right)}}\right\}_{t\geq 1}$, eqn. ˜21 shows how we construct the response and predictor matrix Y and X from $\mathopen{}\mathclose{{\left\{Z^{t}}}\right\}_{t\geq 1}$ for estimating $\Phi$, the matrix formed by stacking $d$ VAR transistion matrices. $\mathsf{Autotune}$ VAR(d) performs $p$ separate $\mathsf{autotune}$ Lasso regression of each column of Y on X to yield $\widehat{\Phi}$.
$\widehat{\Phi}_{\cdot,i}=\underset{\beta\in\mathbb{R}^{dp}}{\operatorname{argmin}}\frac{1}{2n}\mathopen{}\mathclose{{\left\|\text{Y}_{\cdot,i}-\text{X}\beta}}\right\|^{2}+\lambda_{i}\mathopen{}\mathclose{{\left\|{\beta}}}\right\|_{1},\quad\text{ for }i=1,\ldots,p.$ (31)
Separate regularization for each Lasso regression allows $\mathsf{autotune}$ Lasso to adapt to heterogeneity in noise variance across different time series components.
3.2 Noise estimation and sparsity diagnostics with $\mathsf{autotune}$
The intermediate PR and $\widehat{\sigma}^{2}$ of our algorithm show promising results for downstream tasks in high-dimensional regression. In Sec.˜4.3, we show two such examples. First, our $\widehat{\sigma}^{2}$ shows competing accuracy as the state-of-the-art alternatives, and seems to help with high-dimensional inference with bias corrected Lasso. Second, the rank list of PR standard deviations seems to be a useful to diagnose if the underlying model is sparse. We track the improvement in cumulative and adjusted R2 as we build our nested sequence of linear models, and plot these to see any obvious elbow pattern as in scree plot for principal component analysis.
3.3 An illustration of PR properties
Figure 3: Partial residuals’ SD of first 2 coordinate descents (CDs) of $\mathsf{autotune}$ Lasso, which requires 30 iterations of CDs to converge for a high-dimensional setup where $n=80,p=750$. We denote significant predictors (predictors whose true coefficients are non-zero) with “Sig Beta” labeling and green bars, and insignificant predictors are denoted by blue and maroon bars based on whether their coefficients’ estimates are zero/nonzero at the end of that iteration. We also cap the number of insignificant predictors in blue group (their coefficients estimated as zero) to 10 for readability of the plot. Green and blue dashed vertical lines indicates SD of the true residuals and noise SD estimates at the end of the iteration.
We empirically illustrate the behavior of partial residuals in our $\mathsf{autotune}$ Lasso. In Fig.˜3, we simulate the following linear model
$\text{y}_{i}=\text{X}_{i,\cdot}^{\top}\beta_{0}+\epsilon_{i},$ $\quad\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}),\quad\text{for all}\quad i\in\mathopen{}\mathclose{{\left\{1,\dots,n}}\right\}.$ (32)
with $n=80$ observations, $p=750$ covariates, and $\beta_{0}=(1,1,1,1,1,0,0,\dots)\in\real^{p}$, i.e., yi’s depend on only the first 5 covariates. The $\text{X}_{i,\cdot}$’s are independently generated from $p$-dimensional standard normal distribution. Fixing $\text{X}_{i,\cdot}$’s, we generate yi’s multiple times by drawing errors $\epsilon_{i}$’s from $\mathcal{N}(0,\sigma^{2})$ with $\sigma^{2}$ chosen so that Signal-to-Noise Ratio is 1.
We exhibit the histogram of the standard deviations (SD) of all the partial residuals ˜4 obtained from the estimates of $\beta$ at the end of the first two iterations of coordinate descent in autotune Lasso (Algorithm 3) . The predictors are categorized into three groups: Significant predictors i.e., predictors whose true coefficients are non-zero, are depicted in green, and spurious predictors whose estimated coefficients are zero/nonzero are shown in blue and maroon bars, respectively.
In Fig.˜3, since the predictors $\text{X}_{\cdot,j}$ are uncorrelated, the SDs of the partial residuals (PRs) of significant predictors separate from the cluster of PRs of insignificant predictors right from the first iteration. We highlight in that replication, $\mathsf{autotune}$ Lasso took 30 rounds of coordinate descent to converge at the final estimate of $\beta$. In the second iteration, the separation between significant and insignificant predictors becomes wider. Now, we can carve out the support set by correctly identifying the outliers in this histogram via Algorithm 2.
We observe this phenomenon in most of the simulation setups we consider (Sec.˜4.1.1). Note that, we arrive at this by only running one round of CD on $\widehat{\beta}$ at fixed $\lambda_{0}\widehat{\sigma}^{2}$. Hence, inspired by this clustering behavior, we hypothesized that partial residuals would play a crucial role in high-dimensional regression when the sparsity assumption holds.
In general, support set recovery in Lasso is hard, especially in presence of correlated predictors [aos_support_set_recovery], where $\mathsf{autotune}$ can identify some of the insignificant predictors as outliers. But we only need PRs for $\sigma$ estimation, and even if the insignificant predictors which are strongly correlated with significant predictors, replaces the true predictors in the $\mathsf{autotune}$’s estimated support set, then also $\mathsf{autotune}$’s $\widehat{\sigma}$ doesn’t get hampered. We empirically validate this in Sec.˜4.3.1.
4 Numerical Experiments
We demonstrate the performance of $\mathsf{autotune}$ Lasso in three settings: 1) in high-dimensional regression, we compare its runtime, variable selection, estimation of $\beta$, and prediction accuracy against Scaled Lasso and CV, 2) in large VAR models, we compare its runtime, variable selection and estimation accuracy against [kock2025data] and TSCV, 3) we compare the $\widehat{\sigma}^{2}$ of autotune against Scaled Lasso, Organic and Natural Lasso, and illustrate how the PR can be used to design diagnostic plots and check the sparsity assumption.
Note that all our empirical experiments are run in Linux version 5.19.17.
4.1 High-dimensional regression
We closely follow the simulation setup of [hastie2020best], where the data-generating process is completely characterized by 6 parameters: $n$, $p$, $s$ (level of sparsity), $\rho$ (correlation between the predictors), $\mathrm{SNR}$ (Signal to Noise Ratio), and beta-type (pattern of sparsity). We use the following linear model:
$\text{Y}=\text{X}\beta+\epsilon\quad\text{where}\quad\text{Y}\in\real^{n},\quad\text{X}\in\real^{n\times p},\quad\epsilon\sim\mathcal{N}\mathopen{}\mathclose{{\left(0,\sigma^{2}\mathrm{I}}}\right).$ (33)
[…]
Following [hastie2020best], the default value for SNR is set to 2, for simulating correlated predictors, we will primarily use $\rho=0.35$ and we consider five different schemes for generating $\beta\in\mathbb{R}^{p}$:
  1. (B1)
    Beta-type 1: $\beta$ has $s$ entries equal to 1, placed equidistantly between indices 1 and $p$, with all other entries set to 0.
  2. (B2)
    Beta-type 2: The first $s$ entries of $\beta$ are equal to 1, and the remaining entries are 0.
  3. (B3)
    Beta-type 3: The first $s$ entries of $\beta$ form an arithmetic progression series from 10 down to 0.5, while the rest are set to 0.
  4. (B4)
    Beta-type 4: (Approximate Sparsity) The first $s$ entries of $\beta$ are equal to 1, and the remaining entries decay exponentially, specifically $\beta_{i}=0.5^{i-s}$ for $i=s+1,\dots,p$.
  5. (B5)
    Beta-type 5: $\beta$ has $s$ nonzero entries in an arithmetic progression series from 10 down to 0.5 and placed roughly equally spaced across the index range $1$ to $p$, with all other entries set to 0.
Benchmarks:
The main benchmarks CV Lasso and Scaled Lasso are implemented via $\mathsf{cv.glmnet}$ function (from $\mathsf{glmnet}$ package [glmnet]) and $\mathsf{scalreg}$ function (from $\mathsf{scalreg}$ package [scalreg]) respectively. For IC-based tuning, we first fit the whole Lasso path using the $\mathsf{glmnet}$ R package and then choose the final $\lambda$ via AIC or BIC defined in Sec.˜2.
Performance Metrics:
Let $\beta=\mathopen{}\mathclose{{\left(\beta_{1},\dots,\beta_{p}}}\right)$ be the true coefficient, $\widehat{\beta}=\mathopen{}\mathclose{{\left(\widehat{\beta}_{1},\dots,\widehat{\beta}_{p}}}\right)$ be the algorithm’s estimated coefficient vector, and $\text{Y}_{\mathrm{test}}|\text{X}_{\mathrm{test}}\sim\mathcal{N}\mathopen{}\mathclose{{\left(\text{X}_{\mathrm{test}}\beta,\sigma^{2}\mathbf{I}}}\right)$ where the rows of $\text{X}_{\mathrm{test}}$ are drawn from $\mathcal{N}_{p}(0,\Sigma)$, i.e., rows of X and $\text{X}_{\mathrm{test}}$ are i.i.d.. We quantify estimation and prediction error using the Relative Mean Square Error(RMSE) and Relative Test Error(RTE), respectively. RTE measures the ratio of expected prediction error using the algorithm’s estimates $\widehat{\beta}$ to that of the oracle, which knows the true $\beta$.
$\sqrt{\mathrm{RMSE}}=\frac{\mathopen{}\mathclose{{\left\|{\widehat{\beta}-\beta}}}\right\|_{2}}{\mathopen{}\mathclose{{\left\|{\beta}}}\right\|_{2}}=\frac{\sqrt{\sum_{i=1}^{p}\mathopen{}\mathclose{{\left(\widehat{\beta}_{i}-\beta_{i}}}\right)^{2}}}{\sqrt{\sum_{i=1}^{p}\beta_{i}^{2}}},\quad\mathrm{RTE}=\frac{\mathopen{}\mathclose{{\left(\widehat{\beta}-\beta}}\right)^{\top}\Sigma\mathopen{}\mathclose{{\left(\widehat{\beta}-\beta}}\right)+\sigma^{2}}{\sigma^{2}}=\frac{\mathbb{E}\mathopen{}\mathclose{{\left\|{\text{Y}_{\mathrm{test}}-\text{X}_{\mathrm{test}}\widehat{\beta}}}}\right\|_{2}^{2}}{\mathbb{E}\mathopen{}\mathclose{{\left\|{\text{Y}_{\mathrm{test}}-\text{X}_{\mathrm{test}}\beta}}}\right\|_{2}^{2}}.$ (34)
Note that if $\widehat{\beta}=0$, i.e., the algorithm fits a null model on the data, then RMSE =1=1 and RTE $\mathrm{SNR}+1$, while for perfect fit, i.e., $\widehat{\beta}=\beta$, RMSE =0=0 and RTE =1=1. So, a good frame of reference for RMSE is $(0,1)$, while RTE varies in the range $(1,1+\mathrm{SNR})$. For variable selection, we use Area Under the Receiver Operating Characteristic curve(AUROC), a measure of how well an algortihm recovers the sparsity pattern of $\beta$.
Interpretation of SNR:
Our definition of SNR is in line with the SNR in [hastie2020best]. They gave a natural interpretation of the SNR in terms of a quantity called the Proportion of Variance Explained(PVE).
$\mathrm{PVE}(\widehat{\beta})=1-\frac{\mathbb{E}\|\text{Y}_{\mathrm{test}}-\text{X}_{\mathrm{test}}\widehat{\beta}\|_{2}^{2}}{\mathrm{Var}(\text{Y}_{\mathrm{test}})}=1-\frac{\mathopen{}\mathclose{{\left(\widehat{\beta}-\beta}}\right)^{\top}\Sigma\mathopen{}\mathclose{{\left(\widehat{\beta}-\beta}}\right)+\sigma^{2}}{\beta^{\top}\Sigma\beta+\sigma^{2}}$ (35)
Notice that the PVE of the optimal prediction $\widehat{\beta}=\beta$ is PVE($\beta$) $\frac{\mathrm{SNR}}{\mathrm{SNR}+1}$. So, SNR sets the best possible PVE that can be achieved by a linear model, hence it serves as a good indicator in the simulations of how hard is the regression problem in terms of prediction. For instance, SNR of 1 means that linear model can explain atmost 50% of the total variation in the response variable Y.
Sample size, dimension and sparsity:
We will benchmark $\mathsf{autotune}$ against alternative methods in following three regimes:
Dimension $n$ $p$ $s$
Low 80 50 5
Moderate 80 150 5
High 80 750 5
Table 1: Configurations
4.1.1 Effect of Autotune’s tuning in estimation accuracy
[…]
$\mathsf{autotune}$ picks the $\lambda$ closest to the RMSE-optimal $\lambda$ of Lasso solution path in both moderate and high dimensions, yielding $\widehat{\beta}$ which is almost as good as the best $\beta$ estimate one can get from the Lasso solution path after tuning via CV/AIC/BIC. On the other hand, AIC and BIC completely fail to tune the $\lambda$ in the $p>n$ regime. CV(min) and CV(1se) both overfits in moderate dimensional dataset and underfits in high dimensional dataset.
Moderate dimension, $\rho=0.35$, beta-type 1, SNR=2=2 High dimension, $\rho=0.35$, beta-type 1, SNR=2=2
Figure 4: Regularization path of autotune Lasso (blue dots) against Lasso evaluated over its default lambda grid (red dots), vertical lines indicate the final $\lambda$ selected by different procedures. Round blue dots indicate the solution path of $\mathsf{autotune}$ when $\sigma$ is being estimated (hence $\lambda=\lambda_{0}\widehat{\sigma}^{2}$ is changing) and starred blue dot display the final $\widehat{\beta}$ of $\mathsf{autotune}$.
4.1.2 Estimation, prediction and variable selection accuracy
We plot the RMSE, RTE, and AUROC of estimates $\widehat{\beta}$ given by $\mathsf{autotune}$, CV Lasso, and Scaled Lasso against a SNR grid of 20 evenly spaced values on a log scale from 0.66 to 4 in moderate and high dimensions (Tab.˜1). We report the mean of these three metrics across 200 replications along with 1 SE bars. For each setting, we considered all sparsity patterns (all beta-types) and three levels of correlation between predictors, namely $\rho\in\mathopen{}\mathclose{{\left\{0,0.35,0.7}}\right\}$ (same as in [hastie2020best]). A full suite of our simulations is provided in the Sec.˜A.2.
Our primary focus in the paper is the performance of Lasso in $n\ll p$ regime, which is realized in the high-dimensional configuration of Tab.˜1. We noticed that for SNRs below 0.66 (max PVE drops to 40%), all the methods, including autotune, are not doing any meaningful estimation (RMSE nearly 1), hence our SNR grid starts from 0.66. On the other side, we observe that all methods achieve the perfect AUROC of 1 in a large range of setups in high dimensions at SNR =4=4, so we set the maximum SNR to 4.
Results:
We present a slice of the empirical comparisons in Fig.˜5 and Fig.˜6 to keep this section concise and defer the rest in Appendix (Sec.˜A.2). In this setting, $\rho=0.35$ and true coefficients follow the pattern of either beta-type 1 or beta-type 2. In beta-type 2 correlated design, the significant predictors are correlated with one another while for beta-type 1 there is practically no correlation between the significant predictors.
$\mathsf{autotune}$ uniformly shows better variable selection in terms of AUROC than CV and Scaled Lasso across the full range of setups we considered in moderate and high-dimensional settings. We also observe that in almost all settings where a benchmark method shows significant improvement in estimation or prediction accuracy over $\mathsf{autotune}$, that benchmark has an AUROC of 1 in that setup, indicating perfect model selection.
$\mathsf{autotune}$ shows significantly smaller prediction error than the benchmarks across most of the settings considered. In fact, Scaled Lasso struggles a lot in terms of RTE across all settings of high dimensional setup.
In estimation, there are 2 different scenarios. For Beta-type 1, $\mathsf{autotune}$ shows significant improvement in estimation (RMSE) over the benchmarks for all SNRs in both moderate and high-dimensional setups. Only exception was SNRs >3>3 in high dimensional setup, CV Lasso estimates better than $\mathsf{autotune}$. Scaled Lasso struggled to give any competitive performance for beta-type 1. Introducing correlation within significant predictors (Beta-type 2) reduced the RMSE of all algorithms (also observed in [reid2016study]). Scaled Lasso’s performance in estimation now matches that of $\mathsf{autotune}$ for SNRs below a critical value, after which scaled Lasso significantly becomes the best estimator. These critical values are different for moderate and high dimensions, and interestingly, scaled Lasso has AUROC of 1 for all SNRs at or beyond this critical value.
   High Dimensional setup: $n=80,\hskip 2.84544ptp=750,\hskip 2.84544pts=5$, correlation: $\rho=0.35$
Top figure: Beta-type 1 and bottom figure: Beta-type 2
Figure 5: RMSE, RTE, and AUROC of $\mathsf{autotune}$, CV, and Scaled Lasso plotted as a function of SNR for high-dimensional setup.
   Moderate Dimensional setup: $n=80,\hskip 2.84544ptp=150,\hskip 2.84544pts=5$, correlation: $\rho=0.35$
Top figure: Beta-type 1 and bottom figure: Beta-type 2
Figure 6: RMSE, RTE, and AUROC of $\mathsf{autotune}$, CV, and Scaled Lasso plotted as a function of SNR for high-dimensional setup.
4.1.3 Runtime comparsion
We assess the speed of the algorithms through their mean runtime (in milliseconds) for estimating regression coefficients $\beta$ across 100 replications. Standard error bars across replications are also displayed in each figure.
We perform the experiment on datasets generated by fixing $\rho=0.35,\mathrm{SNR}=1$, Beta-type 2 and focusing on 2 settings i) $n=80,s=5$ and vary $p$ from 50 to 750 and ii) $n=200,s=5$ and vary $p$ from 100 to 1000.
(a) $n=80,s=5$, $\rho=0.35$, $p$ varies 50 to 750
(b) $n=200,s=15$, $\rho=0.35$, $p$ varies 100 to 1000
Figure 7: Runtime of $\mathsf{autotune}$, scaled Lasso and CV-Lasso with increasing $p$ for fixed $n,s$. $\mathsf{autotune}$ runs 1510015-100 times faster, with higher gain in speed in high-dimensional settings.
Table 2: Table of mean runtimes (with SD in the brackets) for Fig.˜7(a). $n=80,s=5,$ $\rho=0.35,\mathrm{SNR}=1$, Beta-type 2 and $p$ varies from 50 to 750.
Algorithm 50 64 80 100 120 160
Autotune 1.0 (0.1) 1.0 (0.1) 1.0 (0.1) 1.2 (0.1) 1.5 (0.1) 1.9 (0.6)
CV Lasso 85.1 (12.9) 134.5 (11.5) 122.3 (9.9) 105.9 (9.4) 126.4 (10.3) 140.1 (10.7)
Scaled Lasso 25.8 (3.6) 28.7 (1.6) 53.9 (4.6) 50.1 (3.0) 70.0 (4.9) 72.8 (5.8)
Algorithm 200 240 320 400 560 750
Autotune 2.4 (0.2) 2.6 (0.5) 3.9 (0.3) 6.5 (0.9) 9.9 (0.9) 11.1 (1.4)
CVLasso 136.3 (12.2) 163.2 (16.1) 180.8 (15.5) 200.1 (17.8) 152.4 (15.8) 165.2 (14.9)
Scaled Lasso 75.8 (6.1) 94.8 (6.6) 100.8 (8.3) 105.9 (10.4) 162.6 (19.7) 200.4 (19.1)
In Fig.˜7, we plot the mean runtimes against increasing number of predictiors $(p)$ and the values are given in Tab.˜2 and 5. $\mathsf{autotune}$ shows 15-112 times reduction in runtime as compared to CV Lasso and over 26-107 times as compared to Scaled Lasso.
4.2 Autotune VAR-Lasso
We compare $\mathsf{autotune}$ VAR against VAR Lasso tuned via TSCV, AIC, BIC, and [kock2025data] in terms of estimation, variable selection and runtime. From here on, we refer to [kock2025data] as “KPS Lasso”.
We reuse the metrics employed in the previous subsection 4.1.2, except RTE as in time series, as forecasting is interesting to econometricians and scientists while in-sample prediction is not. In variable selection, $\mathsf{autotune}$ gave AUROC of nearly 1 in all the simulation setups we considered. Hence, we use another variable selection criterion for comparison, Matthews Correlation coefficient(MCC)
$\text{MCC}=\frac{\text{TP.TN}-\text{FP.FN}}{\sqrt{(\text{TP+FP})(\text{TP+FN})(\text{TN+FN})(\text{TN+FP})}}$ (36)
where TP (and FP) is the number of true positives (and false positives) i.e., number of components of $A$ that are estimated as nonzero and are actually nonzero (zero) in the DGP. Similarly, let TN (and FN) is the number of true negatives (and false negatives). For the binary classification problem of whether $A_{i,j}=0$ or not for all $i,j$, then MCC is the correlation between the actual and estimated binary labels. So, MCC is between 1 and -1, with 1 meaning perfect classification, 0 means the classification is no better than random guessing and -1 means perfect disagreement between true and estimated labels.
[…]
We focus on the cases $n=100$ and $200$ and we fix SNRs $\nu_{j}$ of all the $p$ time series to 2.5 for the experiments in Fig.˜8, Fig.˜9, and Fig.˜10 to keep this section concise. We depict the case of heterogenous SNRs in Fig.˜30 and Fig.˜31, where half of the $p$ time series get SNR 5, other half gets SNR 0.5, and all the other parameters of the simulation are kept same. Dimension of parameter space for both DGPs is $p^{2}$, so for $n=200$, $p=15$ gives a moderate dimensional setup and $p\geq 20$ a high dimensional setup. Our main benchmark is KPS Lasso [kock2025data], which also offers an option of post-Lasso in VAR that reduces the shrinkage bias of Lasso estimator by running “regression in two stages”. Since we are reporting the results of Lasso without any post-Lasso bias correction, to keep the comparison with $\mathsf{autotune}$ VAR fair, we use the original VAR Lasso proposed in [Algo 1, kock2025data] as KPS Lasso, instead of their post-Lasso version. We also use VAR Lasso where tuning is done via AIC, BIC and TSCV separately for each column of $\Phi$ ˜22.
For estimation, $\mathsf{autotune}$’s RMSE remains fairly constant as $p$ increases, whereas RMSE of all other methods increase rapidly with dimension. This pattern becomes more prominent for smaller sample size ($n=100)$. TSCV, AIC and BIC often have RMSE >1, i.e., their estimates are worse than a null model $\widehat{A}=0_{p\times p}$. We highlight that even under such difficult high-dimensional setup, $\mathsf{autotune}$’s RMSE doesn’t explode. $\mathsf{autotune}$ uniformly dominates KPS Lasso in terms of estimation accuracy. Additionally, in $2\times 2$ block VAR DGP, KPS Lasso almost gives a null model fit for $n=100,p\geq 10$ as its RMSE becomes nearly 1, and even for $n=200$, KPS Lasso gives RMSE >0.75>0.75 for $p\geq 16$.
$\mathsf{autotune}$ outperforms all the benchmarks in terms of AUROC and provides a perfect AUROC of 1 across all the setups considered in Fig.˜8 and Fig.˜9. In diagonal VAR DGP, KPS shows significantly better MCC $\mathsf{autotune}$ VAR in $n=200$ setup. In fact, KPS has the perfect MCC of 1 for $p\leq 20$. But with smaller sample size of $n=100$, MCC of KPS quickly deteriotes with increasing dimensionality of estimand $A$ and $\mathsf{autotune}$ starts significantly dominating KPS in high-dimensional regime $p^{2}>n=100$. KPS’s MCC gets significantly worse than $\mathsf{autotune}$ in $2\times 2$ block VAR DGP (Fig.˜9), it even drops below 0.2 for $p\geq 16$ in $n=100$ setup whereas $\mathsf{autotune}$ almost maintains same levels of MCC as in Diagonal VAR setup. VAR Lasso tuned by CV/AIC/BIC have much lower MCC compared to $\mathsf{autotune}$, infact their MCC drops below 0.2 for $p\geq 20$.
In simulations with heterogenous SNRs (Fig.˜30, 31), the main conclusions remain same. Heterogenous SNRs further degrade the performance of VAR Lasso with TSCV, AIC, or BIC tuning in terms of RMSE and AUROC in low sample size setup of $n=100$.
   Diagnonal VAR(1) setup: SNR = 2.5
Top figure: Time Series length 100 and bottom figure: Time Series length 200
Figure 8: RMSE, AUROC, and Matthew Correlation Coefficient plotted as functions of number of time series $(p)$ for $n=100$ (top) and 200200 (bottom) in DGP(i).
2x2 Block VAR(1) setup: SNR = 2.5
Top figure: Time Series length 100 and bottom figure: Time Series length 200
Figure 9: RMSE, AUROC, and Matthew Correlation Coefficient plotted as functions of number of time series $(p)$ for $n=100$ (top) and 200200 (bottom) in DGP(ii).
(a) Diagonal VAR, $n=100$
(b) Diagonal VAR, $n=200$
(c) Block VAR, $n=100$
(d) Block VAR, $n=200$
Figure 10: Comparison of runtimes of Autotune VAR Lasso and benchmarks in two different VAR models with SNR = 2.5
In Fig.˜10, we plot the mean runtimes (with 1 SE bars) against increasing number of time series $p$ for VAR Lasso tuned with Autotune, KPS Lasso, TSCV, and AIC under the two DGPs mentioned above, BIC gave identical runtimes as AIC so we dropped it. For diagonal VAR, $\mathsf{autotune}$ is 58 - 169 times faster than TSCV and over 30 - 66 times faster than KPS Lasso. For Block VAR, $\mathsf{autotune}$ is 63 - 175 times faster TSCV and 26 - 84 times faster than KPS Lasso. AIC Lasso’s runtime was always in between KPS and TSCV VAR Lasso.
Overall, this experiment reinforces our claim that $\mathsf{autotune}$ VAR not only ameliorates running time of VAR but also improves VAR estimation in high dimensions and variable selection over KPS Lasso and popular prior works of VAR-Lasso; VAR Lasso tuned via CV, AIC, and BIC.
4.3 Noise estimation and sparsity diagnostics with $\mathsf{autotune}$
While our main objective in this paper is to design a fast and accurate tuning strategy for Lasso, some intermediate products of our algorithm show promising empirical results in statistical tasks relevant in high-dimensional regression. Here we illustrate some preliminary results for two such tasks, estimation of noise variance $\sigma^{2}$, and a visual diagnostic like screeplot to check the sparsity assumption.
4.3.1 Noise scale estimation for high-dimensional inference
We empirically contrast $\sigma$ estimation of $\mathsf{autotune}$ against Scaled Lasso, Organic Lasso and Natural Lasso [yu2019estimating]. We also investigate the quality of estimated $\sigma$ of different algorithms by using it in inference of high-dimensional regression coefficients via $\mathsf{hdi}$ R package [dezeure2015high]. We report the sensitivity of $\sigma$ estimation to chosen $\alpha$ in the $\mathsf{autotune}$’s sequential F-test (see Sigma-Update) in the Sec.˜A.5.
Performance Metric:
We at first compute the difference between algorithm’s estimated $\sigma^{2}$ and the best empirical estimate of noise level, $\sigma^{2}_{\mathrm{empirical}}=\mathrm{Var}(\text{Y}-\text{X}\beta_{\mathrm{true}})$ i.e., variance of the residuals obtained via regression of Y on X using true coefficients. Then we plot the boxplots of the difference $\widehat{\sigma}^{2}-\sigma^{2}_{\mathrm{empirical}}$ across 100 replications in Fig.˜11.
For inference, we use De-Sparsified Lasso [dezeure2015high] from $\mathsf{hdi}$ R package. It computes the Lasso coefficients for the dataset (Y,X), debiases them and then it needs an estimate of $\sigma$ to compute the p-values for two-sided testing of $\mathcal{H}_{0,j}:\beta_{j}=0$ for $j\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$. Over there, $\widehat{\sigma}$ of $\mathsf{autotune}$ and other benchmarks are supplied and their inference results are compared. We set significance level to 0.05 and if p-value of $j^{th}$ covariate is lesser than 0.05, $\mathcal{H}_{0,j}$ is rejected i.e., it is flagged as a significant predictor and vice versa. Let $\mathcal{S}=\mathopen{}\mathclose{{\left\{j\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}:\beta_{j}\neq 0}}\right\}$ be the true support set, then $s=|\mathcal{S}|$ and the false positive rate (FPR) and power becomes
$\mathrm{FPR}$ $=\sum_{j\in\mathcal{S}^{C}}\mathbb{P}\mathopen{}\mathclose{{\left[\mathcal{H}_{0,j}\text{ was rejected }}}\right]/\mathopen{}\mathclose{{\left(p-s}}\right),~~\mathrm{Power}=\sum_{j\in\mathcal{S}}{\mathbb{P}}\mathopen{}\mathclose{{\left[\mathcal{H}_{0,j}\text{ was rejected }}}\right]/s.$ (37)
We plot the empirical estimates of these metrics across 100 replications in Fig.˜12, and add a horizontal line $y=0.05$ in the FPR diagram for connecting the empirical FPRs with the significance level used.
Benchmarks:
Along with Scaled Lasso, we add the Organic and Natural Lasso as [yu2019estimating] focuses on the estimation of error variance in high dimensional regression and are publicly available in the R package $\mathsf{natural}$. Moreover, [yu2019estimating] found organic Lasso to be uniformly better than the CV Lasso based $\widehat{\sigma}$ [reid2016study], hence we skip CV Lasso.
For better understanding of the difficulty in the experiments, we add an oracle method. It has the knowledge of true support set(predictors whose coefficients are nonzero), hence it performs least squares regression of $Y$ on the true support set and report its MSE as the $\widehat{\sigma}^{2}$.
Lastly, in the inference experiment, we drop the results of Organic and Natural Lasso as their $\sigma$ estimation wasn’t competitive in previous experiment.
Experiment DGP:
We focus on the moderate and high-dimensional setups of Tab.˜1, beta types 1˜(B1) and 2˜(B2) and fix the SNR level at 2. We use $\rho=0.35$ to simulate correlated predictors. For tuning in the benchmark methods, the corresponding default settings of their R packages were used.
Results:
We observe in Fig.˜11 that $\mathsf{autotune}$ estimates the $\sigma$ better than all the non-oracle benchmarks. Infact, paired t-test comparisons between $\widehat{\sigma}$’s of $\mathsf{autotune}$ and the benchmarks across 100 replications reveal that the gains in $\sigma$ estimation from $\mathsf{autotune}$ are significant at 0.05 level. The only exception is the setup where the significant predictors are correlated with one another; Beta-type 2 with $\rho=0.35$ where Scaled Lasso matches the performance of $\mathsf{autotune}$. 95% CI of paired t-test comparison between $\mathsf{autotune}$ and Scaled Lasso had 0 in it. Comparison of Fig.˜11(c) and Fig.˜11(d) empirically suggests correlation within the significant predictors hampers the $\widehat{\sigma}$ of $\mathsf{autotune}$ and improves that of Scaled Lasso, but same phenomenon doesn’t happen when significant predictors have correlation primarily with insignificant ones (compare Fig.˜11(a) and Fig.˜11(b)). This observation is in accordance with the comparison of $\mathsf{autotune}$ and Scaled Lasso in terms of RMSE and RTE in Fig.˜5.
The false positive rate(FPR) of all the algorithms was below the significance level of 0.05 across all 100 replications except one. We display side-by-side violin plots of the power of different algorithms across 100 replications, as violin plots give a better comparison of algorithms in terms of power as compared to a jittered scatter plot. The violin plot of $\mathsf{autotune}$’s $\widehat{\sigma}$ closely matches that of oracle’s $\widehat{\sigma}$ across all the setups considered. On the other hand, Scaled Lasso’s $\widehat{\sigma}$ shows suboptimal power everywhere, even in the setup where significant predictors are correlated.
These two experiments support our claim that the partial residuals-driven $\sigma$ estimation mechanism of our $\mathsf{autotune}$ is superior and also exhibits the maximum power empirically possible via $\widehat{\sigma}$ in Lasso inference while controlling the type-1 error rate below the specified level.
(a) Beta-type 1, $\rho=0$
(b) Beta-type 1, $\rho=0.35$
(c) Beta-type 2, $\rho=0$
(d) Beta-type 2, $\rho=0.35$
Figure 11: Comparison of noise variance estimation in the high-dimensional setup
(a) Beta-type 1, $\rho=0$
(b) Beta-type 1, $\rho=0.35$
(c) Beta-type 2, $\rho=0$
(d) Beta-type 2, $\rho=0.35$
Figure 12: Inference of regression coefficients using the estimated $\widehat{\sigma}$ of different algorithms in high dimensional setup
4.3.2 Visual diagnostics for checking sparsity assumption
We potray the cumulative and adjusted $R^{2}$ of the least squares model formed by incrementally adding predictors to the linear model in the rank order given by $\mathsf{autotune}$ Lasso’s partial residuals. This means in Fig.˜13, a $(s,t)$ point denotes the linear model formed by first $s$ ranked predictors has a cumulative/adjusted $R^{2}$ value of $t$.
We demonstrate the diagnostics for a high dimensional setup of $n=400,p=2000$, SNR =2=2, and predictors are uncorrelated with each other, i.e., $\rho=0$. $s$ is varied between 5 and 10 for the sparse case, 350 for the non-sparse case, and 30 for the semi-sparse case(where it becomes unclear from diagnostics whether sparsity holds or not). We report additional information like the number of iterations of CD $\mathsf{autotune}$ runs to arrive at final $\widehat{\beta}_{\mathsf{autotune}}$, $\mathopen{}\mathclose{{\left\|{\widehat{\beta}_{\mathsf{autotune}}}}}\right\|_{0}$, and estimated support set size (or number of outliers detected by F-tests in the final $\sigma$ update step) at the top of the $R^{2}$ plots.
In Fig.˜13, blue dots refer to predictors with nonzero $\widehat{\beta}$ and red dots means predictors with zero $\widehat{\beta}$. We plot the $R^{2}$ curves till the cumulative $R^{2}$ reaches 1.
The philosophy of our diagnostics is the same as determining the number of principal components from a scree plot. Scientists look for an “elbow” in the scree plot and hope that the principal components appearing before the elbow will give an accurate, sparse (low-dimensional) representation of their high-dimensional data. Similarly, we also try to locate an elbow, a point which is preceded by blue dots with visibly big contributions to the cumulative/adjusted $R^{2}$ and followed by a tight cluster of blue/red dots with marginal contribution to the $R^{2}$. Additionally, the slope of the points will drastically change before and after the elbow spot.
The biggest indicator of sparsity is that the blue dots before the elbow should depict a steep rise in $R^{2}$. Look at Fig.˜13(a), and Fig.˜13(b), where sparsity holds with $s\leq 20$. Sometimes, it’s not clear from the diagnostics whether sparsity holds or not. Consider Fig.˜13(c), where the elbow spot seems to be at the $25^{th}$ predictor. It depicts both the steep rise and slow increase before and after the elbow spot, respectively, but there is not a big change in the slope of the points around the elbow. In Fig.˜13(c), $s$ is 30 and RMSE of $\widehat{\beta}_{\mathsf{autotune}}$ is 0.55. In such experiments where $\mathsf{autotune}$ doesn’t fit well (high RMSE), our diagnostics may not give strong evidence of sparsity.
When sparsity holds, one can gauge the number of important predictors by counting blue dots with visible jumps in the $R^{2}$ plots, just like using scree plot in PCA.
If we are unable to find such an elbow(cumulative R2 plot shows smooth increase) or the elbow spot corresponds to a low cumulative/adjusted R2 (say below 0.5), then we recommend that sparsity doesn’t hold in that dataset. We observe all these negative signs in Fig.˜13(d), which is indeed a non-sparse setup with $s=350$.
Our diagnostics are up to the subjectivity of the user for validating sparsity in Lasso, just like using residual plots to assess whether least square regression gave a good fit to the data. We also highlight that this is the first attempt in the growing Lasso literature to propose an approach to empirically validate the sparsity in high-dimensional data.
(a) $s=5$
(b) $s=10$
(c) $s=30$
(d) $s=350$
Figure 13: Behavior of our sparsity diagnostics for varying levels of sparsity (different values of $s$) when $n=400$ and $p=2000$
5 Financial Application
We implement $\mathsf{autotune}$ on a real-life financial dataset, $\mathsf{sp500}$ from the R package $\mathsf{scalreg}$ [scalreg]. We highlight that $\mathsf{scalreg}$ is the R package of Scaled Lasso, making its provided dataset $\mathsf{sp500}$ a reasonable place for benchmarking $\mathsf{autotune}$ and $\mathsf{sp500}$ also has been used in the prior literature [Sec. 6, wu2020survey] for comparing different Lasso algorithms. The dataset $\mathsf{sp500}$ contains close-of-day percentage change data for $p=492$ out of Standard and Poor 500 stocks and Dow Jones Industrial Average (DJIA) index across $252$ days of the year 2008 (also used in the examples shown in the vignette of $\mathsf{scalreg}$ package). So, one observation in this dataset will be the percentage change on a particular day for 492 stocks and the DJIA index. Our goal is modeling the DJIA using those 492 stocks.
We take the first $n$ observations in our training dataset, and assign the rest of the observations to the test dataset. We vary $n\in\mathopen{}\mathclose{{\left\{150,160,170,180,190,200,210}}\right\}$ and then compare $\mathsf{autotune}$ against CV Lasso, Scaled Lasso, and Lasso with AIC/BIC tuning in terms of Relative Test Error (RTE), Multiple $\text{R}^{2}$ on test data and sparsity of the final model selected, following the comparison metrics used in [wu2020survey]. Unlike Sec.˜4, here we neither know the true regression coefficients nor the true DGP, so we redefine RTE $\mathopen{}\mathclose{{\left\|{\text{Y}_{\mathrm{test}}-\text{X}_{\mathrm{test}}\widehat{\beta}}}}\right\|_{2}^{2}\big/\mathopen{}\mathclose{{\left\|{\text{Y}_{\mathrm{test}}}}}\right\|_{2}^{2}$. Note that with this definition, we have OOS $^{2}=1-R$TE.
We plot the RTE of the solution path taken by the Lasso using the R function $\mathsf{cv.glmnet}$ (orange) and $\mathsf{autotune}$ (blue) on the test data along with normalized CV errors on the training dataset (red) and demarcate the finally chosen $\lambda$’s with labeled vertical lines in Fig.˜14(a), and 34. We choose this splitting mechanism over others as it ensures test error curve has a “U-shape”.
To further validate $\mathsf{autotune}$’s gains in performance, we generate $B=100$ bootstrap samples of our training set, fit $\mathsf{autotune}$ and benchmarks on each sample, and then evaluate the trained model’s Multiple $\text{R}^{2}$ on the test data, i.e. Out-of-Sample $\text{R}^{2}$ (OOS $\text{R}^{2}$). We report these using side-by-side boxplots in Fig.˜14(b), and 35. AIC and BIC Lasso’s results were omitted as they were not competitive.
In Fig.˜14(a), $\mathsf{autotune}$ quickly reaches it final $\lambda$ within 3 iterations for $n=190$, and $\mathsf{autotune}$’s fitted model also has the lowest RTE. The red training curve shows monotone increase in CV error with increase in $\lambda$ for all $n$, suggesting that CV will lead to overfitting (Fig.˜14(a), 34). Tab.˜3 shows $\mathsf{autotune}$ dominates all the benchmarks across all $n$ in terms of RTE.
For bootstrap results, $\mathsf{autotune}$ shows higher median, 25th, and 75th quantile OOS $\text{R}^{2}$ than all its benchmarks for $n=190$ (Fig.˜14(b)). Bootstrapped OOS R2 gives the same results for other choices of $n$ (see Fig.˜35).
$n$ $\mathsf{autotune}$ Scaled CV(min) CV(1se) AIC BIC
150 0.749 0.845 0.832 0.819 0.832 0.826
160 0.633 0.694 0.682 0.676 0.682 0.682
170 0.515 0.567 0.561 0.555 0.561 0.555
180 0.343 0.437 0.437 0.427 0.437 0.437
190 0.250 0.341 0.336 0.331 0.336 0.336
200 0.255 0.340 0.335 0.325 0.335 0.335
210 0.332 0.403 0.397 0.379 0.397 0.391
Table 3: RTE of various algorithms for different training dataset sizes $n$.
(a) RTE of different tuning algorithms.
(b) Boxplot of OOS R2 on 100 bootstraped datasets.
Figure 14: 14(a) shows solution paths of Lasso (evaluated by R function $\mathsf{cv.glmnet}$) in orange and $\mathsf{autotune}$ in blue (with endpoint denoted by blue star) for train/test split with $n=190,p=492$. Finally, selected $\lambda$’s are indicated by labeled vertical lines. Fig.˜14(b) gives the boxplot of OOS R2 (=1=1-RTE) of the algorithms trained on 100 bootstraped datasets of the training dataset with $n=190$. The OOS R2 on the original training dataset is indicated by a red diamond point.
We see that $\mathsf{autotune}$ usually gives the sparsest model among all the benchmarks. Scaled Lasso gives the largest estimated model (also observed in [wu2020survey]). All the remaining benchmark methods (excluding CV(1se)) choose the smallest $\lambda$ in the solution path of Lasso (Fig.˜14(a), 34), resulting in overfitted large models.
$n$ $\mathsf{Autotune}$ Scaled CV.min CV.1se AIC BIC
150 48 65 67 59 67 62
160 57 64 61 58 61 61
170 57 68 60 58 60 54
180 62 63 59 55 59 58
190 47 64 60 57 60 60
200 41 64 61 58 61 61
210 57 70 62 61 62 59
Table 4: Number of nonzero estimated regression coefficients by various algorithms while modeling DJIA.
6 Conclusion
We develop $\mathsf{autotune}$, an automatic tuning parameter selection algorithm for Lasso, and show its performance in high-dimensional regression and large VAR model estimation using simulated data. In contrast to existing automatic tuning procedures such as the Scaled Lasso, $\mathsf{autotune}$ makes use of intermediate partial residuals computed during coordinate descent of lasso to form a rank list of important predictors, and use that list to get a bias-corrected estimator of $\sigma$ which guides tuning parameter selection. The algorithm shows promising results in terms of generalization, model selection and runtime. An R package is also made available on Github.
We did not delve into the convergence analysis of this algorithm in this paper. Characterizing the limit point of our iterative algorithm, and understanding its statistical properties will be crucial to gain insight into conditions under which the algorithm fails. We leave these for future work.
Acknowledgements
SB acknowledges partial support from NSF grants DMS-1812128, DMS-2210675, NSF CAREER award DMS-2239102 and NIH award R01GM135926.
Appendix A Appendix
Contents
A.1 Algorithm of Autotune Lasso with active set selection
Input: Predictors xx(normalized), response $y$, Sequential F-test significance level $\alpha$
1
21exInitialize $\widehat{\beta}\leftarrow 0$, $r\leftarrow y$, $\widehat{\sigma}^{2}\leftarrow\mathrm{Var}(y)$, $\lambda_{0}\leftarrow\frac{1}{\mathrm{Var}(y)}\|\frac{x^{\top}y}{2n}\|_{\infty}$, $\mathsf{Predictor.Ranking}\leftarrow 1:p$, $\mathsf{Support.Set}\leftarrow\phi$, sigma.update.flag $\leftarrow\mathsf{TRUE}$
3
41exwhile $\mathrm{error}\geq 10^{-3}$ do
5  $\mathsf{Support.Set}^{\mathsf{(old)}}\leftarrow\mathsf{Support.Set}$, $\widehat{\beta}^{\mathsf{(old)}}\leftarrow\widehat{\beta}$
6 [2pt]
7 1exfor $i=1,\ldots,p$ do
8     $j\leftarrow\mathsf{Predictor.Ranking}[i]$
9    [2pt] $\widehat{\beta}_{j}\leftarrow\mathsf{Soft.Threshold}_{\lambda_{0}\widehat{\sigma}^{2}}\mathopen{}\mathclose{{\left(\frac{1}{n}x_{\cdot,j}^{\top}r+\widehat{\beta}_{j}}}\right)$
10    [2pt] $r\leftarrow r+x_{\cdot,j}\mathopen{}\mathclose{{\left(\widehat{\beta}^{\mathsf{(old)}}_{j}-\widehat{\beta}_{j}}}\right)$
11  end for
12 Calculate partial residuals $r^{(i)}\leftarrow r+x_{\cdot,i}\widehat{\beta}_{i}$
13
[…]
15
[…]
17 [2pt] if $\mathrm{sigma.update.flag}==\mathsf{TRUE}$ then
[…]
19  end if
20 if $\mathsf{Support.Set}\subseteq\mathsf{Support.Set}^{\mathsf{(old)}}$ then
21     sigma.update.flag $\leftarrow\mathsf{FALSE}$
22     $\mathsf{Support.Set}^{\mathsf{(old)}}\leftarrow\phi$
23     if $\mathsf{Active.Selection}$ is $\mathsf{TRUE}$ then
[…]
25     end if
26
27  end if
28 $\mathrm{error}\leftarrow\frac{\|\widehat{\beta}-\widehat{\beta}^{\mathsf{(old)}}\|_{1}}{\max(\|\widehat{\beta}^{\mathsf{(old)}}\|_{1},10^{-8})}$
29 end while
Output: regression coefficients’ estimate $\widehat{\beta}$
Algorithm 4 Active Autotune Lasso – Return estimates of regression coefficients $\beta$.
Input: $x$, $r$, $\widehat{\beta}$, $\mathsf{Predictor.Ranking}$, $\mathsf{Support.Set}$, thresholding parameter $\lambda$
1
21exInitialize $\mathsf{Active.Set}\leftarrow\mathsf{Support.Set}$
3 for $i=\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left|\mathsf{Support.Set}}}\right|+1}}\right),\ldots,p$ do
4  $j\leftarrow\mathsf{Predictor.Ranking}[i]$
5 [2pt] if $\mathopen{}\mathclose{{\left|\frac{1}{n}\langle{x_{\cdot,j}},{r}\rangle}}\right|\geq\lambda$ then
6     Add $x_{j}$ to the back of $\mathsf{Active.Set}$
7  end if
8 else
9     $r\leftarrow r+x_{j}\widehat{\beta}_{j}$
10     $\widehat{\beta}_{j}\leftarrow 0$
11  end if
12
13 end for
[…]
[2pt] Output: $\widehat{\beta}$, $r$, $\mathsf{Predictor.Ranking}$.
Algorithm 5 Active Set Selector – Performs active set screening of predictors.
A.2 Complete suite of comparison plots in variable selection, estimation, and prediction accuracy
A.2.1 High Dimensional Setting: n = 80, p = 750, s = 5
High Dimensional setup: $n=80,\hskip 2.84544ptp=750,\hskip 2.84544pts=5$, Beta-type: 1, Varying levels
of correlation, Top Row: $\rho=0$, Middle Row: $\rho=0.35$ and Bottom Row: $\rho=0.7$
Figure 15: RMSE, RTE, and AUROC of $\mathsf{autotune}$, $\mathsf{autotune.active}$, CV, and Scaled Lasso plotted as a function of SNR for high-dimensional setup.
High Dimensional setup: $n=80,\hskip 2.84544ptp=750,\hskip 2.84544pts=5$, Beta-type: 2, Varying levels
of correlation, Top Row: $\rho=0$, Middle Row: $\rho=0.35$ and Bottom Row: $\rho=0.7$
Figure 16: RMSE, RTE, and AUROC of $\mathsf{autotune}$, $\mathsf{autotune.active}$, CV, and Scaled Lasso plotted as a function of SNR for high-dimensional setup.
High Dimensional setup: $n=80,\hskip 2.84544ptp=750,\hskip 2.84544pts=5$, Beta-type: 3, Varying levels
of correlation, Top Row: $\rho=0$, Middle Row: $\rho=0.35$ and Bottom Row: $\rho=0.7$
Figure 17: RMSE, RTE, and AUROC of $\mathsf{autotune}$, $\mathsf{autotune.active}$, CV, and Scaled Lasso plotted as a function of SNR for high-dimensional setup.
High Dimensional setup: $n=80,\hskip 2.84544ptp=750,\hskip 2.84544pts=5$, Beta-type: 5, Varying levels
of correlation, Top Row: $\rho=0$, Middle Row: $\rho=0.35$ and Bottom Row: $\rho=0.7$
Figure 18: RMSE, RTE, and AUROC of $\mathsf{autotune}$, $\mathsf{autotune.active}$, CV, and Scaled Lasso plotted as a function of SNR for high-dimensional setup.
A.2.2 Moderate Dimensional Setting: n = 80, p = 150, s = 5
Moderate Dimensional setup: $n=80,\hskip 2.84544ptp=150,\hskip 2.84544pts=5$, Beta-type: 1, Varying levels
of correlation, Top Row: $\rho=0$, Middle Row: $\rho=0.35$ and Bottom Row: $\rho=0.7$
Figure 19: RMSE, RTE, and AUROC of $\mathsf{autotune}$, $\mathsf{autotune.active}$, CV, and Scaled Lasso plotted as a function of SNR for moderate-dimensional setup.
Moderate Dimensional setup: $n=80,\hskip 2.84544ptp=150,\hskip 2.84544pts=5$, Beta-type: 2, Varying levels
of correlation, Top Row: $\rho=0$, Middle Row: $\rho=0.35$ and Bottom Row: $\rho=0.7$
Figure 20: RMSE, RTE, and AUROC of $\mathsf{autotune}$, $\mathsf{autotune.active}$, CV, and Scaled Lasso plotted as a function of SNR for moderate-dimensional setup.
Moderate Dimensional setup: $n=80,\hskip 2.84544ptp=150,\hskip 2.84544pts=5$, Beta-type: 3, Varying levels
of correlation, Top Row: $\rho=0$, Middle Row: $\rho=0.35$ and Bottom Row: $\rho=0.7$
Figure 21: RMSE, RTE, and AUROC of $\mathsf{autotune}$, $\mathsf{autotune.active}$, CV, and Scaled Lasso plotted as a function of SNR for moderate-dimensional setup.
Moderate Dimensional setup: $n=80,\hskip 2.84544ptp=150,\hskip 2.84544pts=5$, Beta-type: 5, Varying levels
of correlation, Top Row: $\rho=0$, Middle Row: $\rho=0.35$ and Bottom Row: $\rho=0.7$
Figure 22: RMSE, RTE, and AUROC of $\mathsf{autotune}$, $\mathsf{autotune.active}$, CV, and Scaled Lasso plotted as a function of SNR for moderate-dimensional setup.
A.2.3 Low Dimensional Setting: n = 80, p = 50, s = 5
Low Dimensional setup: $n=80,\hskip 2.84544ptp=50,\hskip 2.84544pts=5$, Beta-type: 1, Varying levels
of correlation, Top Row: $\rho=0$, Middle Row: $\rho=0.35$ and Bottom Row: $\rho=0.7$
Figure 23: RMSE, RTE, and AUROC of $\mathsf{autotune}$, $\mathsf{autotune.active}$, CV, and Scaled Lasso plotted as a function of SNR for low-dimensional setup.
Low Dimensional setup: $n=80,\hskip 2.84544ptp=50,\hskip 2.84544pts=5$, Beta-type: 2, Varying levels
of correlation, Top Row: $\rho=0$, Middle Row: $\rho=0.35$ and Bottom Row: $\rho=0.7$
Figure 24: RMSE, RTE, and AUROC of $\mathsf{autotune}$, $\mathsf{autotune.active}$, CV, and Scaled Lasso plotted as a function of SNR for low-dimensional setup.
Low Dimensional setup: $n=80,\hskip 2.84544ptp=50,\hskip 2.84544pts=5$, Beta-type: 3, Varying levels
of correlation, Top Row: $\rho=0$, Middle Row: $\rho=0.35$ and Bottom Row: $\rho=0.7$
Figure 25: RMSE, RTE, and AUROC of $\mathsf{autotune}$, $\mathsf{autotune.active}$, CV, and Scaled Lasso plotted as a function of SNR for low-dimensional setup.
Low Dimensional setup: $n=80,\hskip 2.84544ptp=50,\hskip 2.84544pts=5$, Beta-type: 5, Varying levels
of correlation, Top Row: $\rho=0$, Middle Row: $\rho=0.35$ and Bottom Row: $\rho=0.7$
Figure 26: RMSE, RTE, and AUROC of $\mathsf{autotune}$, $\mathsf{autotune.active}$, CV, and Scaled Lasso plotted as a function of SNR for low-dimensional setup.
A.2.4 Accuracy when only approximate sparsity holds
High Dimensional setup: $n=80,\hskip 2.84544ptp=750,\hskip 2.84544pts=5$, Beta-type: 4, Varying levels
of correlation, Top Row: $\rho=0$, Middle Row: $\rho=0.35$ and Bottom Row: $\rho=0.7$
Figure 27: RMSE, RTE, and AUROC of $\mathsf{autotune}$, $\mathsf{autotune.active}$, CV, and Scaled Lasso plotted as a function of SNR for high-dimensional setup.
Medium Dimensional setup: $n=80,\hskip 2.84544ptp=150,\hskip 2.84544pts=5$, Beta-type: 4, Varying levels
of correlation, Top Row: $\rho=0$, Middle Row: $\rho=0.35$ and Bottom Row: $\rho=0.7$
Figure 28: RMSE, RTE, and AUROC of $\mathsf{autotune}$, $\mathsf{autotune.active}$, CV, and Scaled Lasso plotted as a function of SNR for medium-dimensional setup.
Low Dimensional setup: $n=80,\hskip 2.84544ptp=150,\hskip 2.84544pts=5$, Beta-type: 4, Varying levels
of correlation, Top Row: $\rho=0$, Middle Row: $\rho=0.35$ and Bottom Row: $\rho=0.7$
Figure 29: RMSE, RTE, and AUROC of $\mathsf{autotune}$, $\mathsf{autotune.active}$, CV, and Scaled Lasso plotted as a function of SNR for medium-dimensional setup.
A.3 Deferred comparison plots of $\mathsf{autotune}$ VAR against its benchmarks under heterogenous SNRs.
   Diagnonal VAR(1) setup: SNR(s) = 5, 0.5
Top figure: Time Series length 100 and bottom figure: Time Series length 200
Figure 30: RMSE, AUROC, and Matthew Correlation Coefficient plotted as functions of number of time series $(p)$ for n=100n=100(top) and 200200(bottom) in DGP(i).
2x2 Block VAR(1) setup: SNR(s) = 5, 0.5
Top figure: Time Series length 100 and bottom figure: Time Series length 200
Figure 31: RMSE, AUROC, and Matthew Correlation Coefficient plotted as functions of number of time series $(p)$ for n=100n=100(top) and 200200(bottom) in DGP(ii).
A.4 Runtime comparison of $\mathsf{Autotune}$ vs $\mathsf{autotune.active}$ Lasso
Table 5: Table of mean runtimes (with SD in the brackets) for Fig.˜7(b). $n=200,s=15,$ $\rho=0.35,\mathrm{SNR}=1$, Beta-type 2 and $p$ varies from 100 to 1000.
Algorithm 100 160 180 200 250 300
Autotune 4.5 (1.9) 4.6 (0.7) 5.6 (1.6) 7.9 (0.5) 7.1 (0.5) 7.5 (0.6)
CVLASSO 140.0 (19.6) 482.2 (45.9) 1203.1 (76.4) 521.5 (32.7) 411.4 (26.8) 434.8 (28.2)
Scaled Lasso 79.9 (31.8) 270.0 (93.6) 332.1 (28.7) 749.7 (51.5) 613.4 (39.5) 664.0 (46.0)
Algorithm 400 500 600 800 1000
Autotune 13.2 (1.1) 15.5 (1.9) 20.5 (3.1) 34.2 (13.0) 34.9 (5.9)
CVLASSO 513.6 (25.3) 303.0 (40.5) 346.4 (27.7) 373.3 (26.3) 422.9 (32.1)
Scaled Lasso 775.6 (51.6) 821.5 (63.7) 962.8 (94.9) 1272.5 (149.7) 1383.9 (100.4)
(a) Beta-type 1, $\rho=0$, SNR = 2
(b) Beta-type 1, $\rho=0.35$, SNR = 2
(c) Beta-type 2, $\rho=0$, SNR = 2
(d) Beta-type 2, $\rho=0.35$, SNR = 2
Figure 32: Runtime of $\mathsf{autotune}$ lasso with its active set version, $\mathsf{autotune.active}$ lasso as a function of increasing $p$ for $n=400,s=5$.
A.5 Sensitivity of $\sigma$ estimation to $\alpha$
Figure 33: Comparison of $\mathsf{Autotune}$ Lasso’s noise variance estimation in the high-dimensional setup among different choices of $\alpha$ in the inner sequential F-test (see Sigma-Update).
A.6 Supplementary plots for Data Analysis Section
(a) train set size $n=150$, test set size 252n=102252-n=102.
(b) train set size $n=170$, test set size 252n=82252-n=82.
(c) train set size $n=180$, test set size 252n=72252-n=72.
(d) train set size $n=200$, test set size 252n=52252-n=52.
(e) train set size $n=210$, test set size 252n=42252-n=42
Figure 34: Solution paths of CV Lasso in orange and $\mathsf{autotune}$ in blue (with endpoint denoted by blue star) for multiple train/test splits. Finally, selected $\lambda$’s are indicated by labeled vertical lines.
(a) train set size $n=150$, test set size 252n=102252-n=102.
(b) train set size $n=170$, test set size 252n=82252-n=82.
(c) train set size $n=180$, test set size 252n=72252-n=72.
(d) train set size $n=200$, test set size 252n=52252-n=52.
(e) train set size $n=210$, test set size 252n=42252-n=42
Figure 35: Boxplot of 100 bootstraped Multiple R2 on test data for different tuning strategies.